### v2

# Model order reduction via Lie groups

Abstract
Lie groups and their actions are ubiquitous in the description of physical systems, and we explore implications in the setting of model order reduction (MOR). We present a novel framework of MOR via Lie groups, called MORLie, in which high-dimensional dynamical systems on manifolds are approximated by low-dimensional dynamical systems on Lie groups. In comparison to other Lie group methods we are able to attack non-equivariant dynamics, which are frequent in practical applications, and we provide new non-intrusive MOR methods based on the presented geometric formulation. We also highlight numerically that MORLie has a lower error bound than the Kolmogorov $N$-width, which limits linear-subspace methods. The method is applied to various examples: 1. MOR of a simplified deforming body modeled by noisy point cloud data following a sheering motion, where MORLie outperforms a naive POD approach in terms of accuracy and dimensionality reduction. 2. Reconstructing liver motion during respiration with data from edge detection in MRI scans, where MORLie reaches performance approaching the state of the art, while reducing the training time from hours on a computing cluster to minutes on a mobile workstation. 3. An analytic example showing that the method of freezing (a previous Lie group method for MOR introduced in [1]) is analytically recovered as a special case, showing the generality of the geometric framework.
1 Introduction
Figure 1: Summary of MORLie. Given a full order model (FOM) defined by a vector field $X$ on a state-manifold $\mathcal{M}$, we construct a reduced order model (ROM) on a lower-dimensional Lie group $G$. A group action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ is used to reconstruct states via $\bar{x}(t)=\Phi(g(t),x_{0})$, where the evolution of $g(t)$ is determined by a reduced vector field $\rho:\mathcal{M}\rightarrow\mathfrak{g}$. The ROM dynamics on $G$ are defined as $\dot{g}={R_{g}}_{*}\rho(\bar{x})$, and induce approximated dynamics on $\mathcal{M}$ that are constrained to group orbits. The goal is to choose $(G,\Phi,\rho)$, such that the reconstructed solutions $\bar{x}(t)$ closely follows the FOM dynamics $\dot{\bar{x}}\approx X(\bar{x})$.
Model order reduction (MOR) is an essential step for the repeated simulation and optimization of distributed and multi-scale engineering systems, such as soft robotic manipulators [2, 3, 4], hybrid reaction-diffusion systems [5, 6], fluids [7] and systems with fluid-structure interactions [8]. These systems are frequently modeled by (parameterized) partial differential equations (PDEs), which yield high dimensional ordinary differential equations (ODEs) upon spatial discretization – often considered as full order models (FOMs). The core idea of MOR is to approximate these FOMs by low-dimensional surrogate models, called reduced order models (ROMs), that can be evaluated with far lower computational complexity.
A big class of MOR techniques begins with large datasets of snapshots, which could be solutions from high-fidelity simulations of FOMs or state measurements from real world experiments, for an interesting range of system parameters and initial conditions. In the following, we denote the high-dimensional state-manifold of the FOM as $\mathcal{M}$, a set of solution snapshots (classically, the solution manifold) as $S\subseteq\mathcal{M}$, and we lump parameters and initial conditions as $\mu\in\mathcal{P}$, with $\mathcal{P}$ a compact set. In linear-subspace methods [9, 10], $\mathcal{M}$ is a Hilbert space, and $S$ is approximated by a subspace $W\subseteq\mathcal{M}$, e.g., spanned by suitable reduced basis elements $\{\varphi_{1},\cdots,\varphi_{N}\},\;\varphi_{i}\in W$ [11]. Introducing time-dependent coefficients $c_{\mu}^{i}(t)\in\mathbb{R}$ indexed by $i$, real solutions $u_{\mu}(t)\in S$ are approximated as
$u_{\mu}(t)\approx\sum_{i=1}^{N}c_{\mu}^{i}(t)\varphi_{i}\,.$ (1)
Linear-subspace methods scale to high-dimensional systems, can preserve physical structure and can capture input-output behavior for control applications [12, 13]. Theoretical error bounds are well-known: the Kolmogorov $N$-width [14] gives, depending on $S$, a lower bound to the approximation error that can be achieved by an $N$-dimensional subspace $W$ of $\mathcal{M}$. However, linear-subspace methods are ill-suited for $S$ with a slowly decaying Kolmogorov $N$-width, e.g., in certain transport dominated problems from fluid dynamics or heat transport, where solutions do not evolve on fixed subspaces [1].
This slow decaying Kolmogorov $N$-width is often referred to as the Kolmogorov barrier, and to break it, nonlinear MOR methods are required [15], for example by allowing also the basis elements to change over time:
$u_{\mu}(t)\approx\sum_{i=1}^{N}c_{\mu}^{i}(t)\varrho_{\mu,i}(t)\,.$ (2)
Research on MOR methods going beyond the Kolmogorov barrier include: adaptive basis methods [16, 17], which recompute the reduced basis over time, polynomial approximation methods [18, 19, 20], and machine learning approaches [21, 22, 23, 19, 24], which are sufficiently complex to express relations of the form (2) but typically incur large computational cost.
A recent differential geometric description of MOR encompasses many of the aforementioned methods in a common description as submanifold methods [25]: given the FOM as a dynamical system on a $\mathcal{M}$, the ROM is a dynamical system on a submanifold (rather than a subspace) $\mathcal{N}$ of $\mathcal{M}$.
In the present article we use a new geometric approach, utilizing the action of Lie groups. Lie groups act on the underlying configuration and state manifolds for a wide variety of physical systems, including rigid bodies, flexible bodies and fluids, where they allow to describe and exploit symmetries of the systems’ kinematics and dynamics [26]. With reference to Figure 1 our approach reduces a dynamical system on a high-dimensional state-manifold $\mathcal{M}$ to a lower-dimensional dynamical system on a Lie group $G$. Let $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ denote an action of $G$ on $\mathcal{M}$, and $u_{\mu,0}\in\mathcal{M}$ the initial condition of an FOM trajectory, then we work towards approximations of the type
$u_{\mu}(t)\approx\Phi(g_{\mu}(t),u_{\mu,0})\,.$ (3)
While submanifold methods for MOR restrict trajectories to lie on a fixed submanifold, we show that (3) instead restricts the dynamics to directions tangent to group orbits. Thus, rather than constraining where states can lie, we constrain how they are allowed to move. In this way, initial conditions are preserved: for any initial state $u_{0}\in\mathcal{M}$, the approximation evolves on the orbit through $u_{0}$, so no projection onto a reduced manifold is required. More formally, our method induces a low-dimensional distribution on the state-space manifold, rather than selecting a submanifold of it.
Early Lie group methods for MOR [1, 27, 28] split dynamics into group and manifold parts (vertical and horizontal components on a principal bundle, respectively) and applied submanifold methods to reduce the latter. The splitting explicitly required equivariance of the underlying dynamical systems, and MOR methods were intrusive, i.e., based on explicit knowledge of the FOM dynamics. Our method does not require equivariance, and allows a novel non-intrusive (i.e., data-driven) approach that exploits the presence of approximate group motions. Recently [29] used a learning approach to find a ROM on a subgroup of $G=\mathrm{Diff}(\mathbb{R}^{d})$ to reduce Burgers equation with discontinuous shocks. Our method works towards a general geometric approach enabling other choices of $G$, in a fashion that is compatible with machine-learning, but does not rely on it.
Concretely, our contributions are:
  1. 1.
    Development of a new MOR procedure via Lie groups, which we name MORLie (Section 4.1).
  2. 2.
    Formalization of a group Kolmogorov $N$-width, providing a lower bound for the approximation error of MORLie (Section 4.2).
  3. 3.
    Development of optimization problems to compute ROMs, both intrusive and non-intrusive, and numerical methods to solve them (Section 5).
  4. 4.
    Multiple example applications: analytically recovering the method of freezing [1] (Section 6), point clouds undergoing sheering motions and liver tracking during respiration (Section 7).
The article is ordered as follows. We conclude the introduction with related literature and notation. Section 2 provides background on Lie groups and their actions on manifolds. Section 3 introduces the main problem, i.e., to describe an MOR method that results in a ROM on a Lie group and to conceptually connect such a method to existing theory on MOR on manifolds. Next, Section 4 formally describes MORLie and introduces the group Kolmogorov $N$-width. Afterwards, Section 5 presents intrusive and non-intrusive optimization methods, and initial ideas for hyperreduction. Section 6 provides applications of MORLie to linear transport and the method of freezing [1]. Section 7 provides numerical example applications and a discussion of the results. A conclusion and outlook are given in Section 8.
1.1 Related literature
The review [13] presents a comparative study of linear-subspace MOR in structural mechanics, describing modal truncation techniques, acceleration methods, Krylov subspace based MOR and balanced truncation. In [30], center manifold reduction as well as the Galerkin method, modal synthesis method and proper orthogonal decomposition (POD) are reviewed, where the latter three are linear-subspace methods. The review [31] focusses on interconnection-based MOR for both linear-subspace cases and nonlinear applications. In [32] various linear-subspace techniques and nonlinear methods are reviewed, also putting a focus on a posteriori error bounds and the Kolmogorov $N$-width of the to-be-reduced systems.
Submanifold methods for MOR are formalized in a differential geometric language in [25]. Most nonlinear approaches to MOR can be seen as particular cases of submanifold methods: transformation-based methods [33, 34, 35, 36] apply linear or nonlinear, time-dependent transformations to a fixed and reduced basis (still allowing for time-varying coefficients). A similar formalism can also be derived by transforming snapshot matrices with nonlinear transformations [37], which can reduce the Kolmogorov $N$-width. By transforming a reduced basis, such methods effectively represent the system on a submanifold of the state-space. The work of [38, 39] uses a shifted POD, where the shift can be seen as a particular choice of Lie group action – we generalize towards accommodating arbitrary Lie group actions, see e.g., Section 6.2. In [40] integration on Lie groups $SO(n),SP(n)$ is used for geometrically exact integration of time-varying reduced bases for Hamiltonian systems. We interpret this as a submanifold method, since the Lie groups are made to act on a Stiefel (sub-)manifold [41] that the time-varying reduced basis of the total system.
Various machine learning approaches fall into the category of submanifold methods for MOR: in [21] deep convolutional autencoders are applied for manifold Galerkin projection and manifold least-squares Petrov-Galerkin projection, the approach overcomes the Kolmogorov $n$-width, outperforming linear subspace methods on Benchmark 1D Burgers equation and chemically reacting flows, and aposteriori error estimates are provided. ROMs for nonlinear parameterized PDEs are learned by separately identifying deep neural nets parameterizing the submanifold and the reduced dynamics, in [42]. Structure-preserving MOR method for Hamiltonian systems is presented in [22], where deep autoencoders are used and symplecticity is enforced softly as an additional cost-function term, strongly improving convergence over approaches that do not preserve structure. The machine learning approach [43] learns a 2D subspectral manifold to represent Couette flow and its bifurcations at low Reynolds numbers. Autoencoders, which naturally parameterize submanifolds, are applied to achieve MOR of compressible and incompressible Navier-Stokes equations for a large array of examples, including flow around a wing for different Mach numbers, in [24]. In [44] reduced-order neural operators on sparse graphs are learned in a physics-informed manner, and in [45] data on fixed point-clouds is used to learn physics from partial information, using a reduced latent representation that parameterizes a submanifold. In [19] a multi-stage neural network based approach for mitigating the Kolmogorov $N$-width is applied to shock-dominated, unsteady flow and hyperreduction is achieved. A submanifold is composed of an initial reduced basis and additional, nonlinear input from the neglected basis elements.
The method of freezing [32] highlighted earlier, splits certain equivariant PDEs into a Lie group and a vector-space part, applying linear-subspace MOR to the latter. It has its background in reduction of equivariant dynamic systems [46, 47, 48], which had early applications to MOR of equivariant PDEs in [28, 27].
Some recent machine learning approaches to fluid- and continuum mechanics do not fit into a submanifold framework, but can be interpreted as instances of the general Lie group methods we present: [23] learn deformation maps for continuum mechanics applications, achieving a reduced-order latent-state representation that is grid-agnostic, allowing hyperresolution during reconstruction of the full state. In [49] a similar approach is applied to both viscid/ inviscid and compressible/incompressible fluid dynamics. Both approaches implicitly learn representations of the Lie group $\mathrm{Diff}(\mathbb{R}^{3})$. In [29] the Lie group $\mathrm{Diff}(\mathbb{R}^{d})$ is explicitly parameterized by integrating a finite set of neural net parameterized vector fields, which are optimized in a non-intrusive fashion for the example of the 2D Burgers equation, where the resulting method is able to represent shocks.
Apart from [29], applications of Lie groups in MOR were restricted to equivariant dynamic systems, and according to our best knowledge no work in MOR explicitly treats approximate equivariance. Within the machine learning community, both equivariant and approximately equivariant systems were investigated in detail: [50] investigates approximately equivariant networks for learning approximately equivariant dynamics, and apply this to learning dynamics of smoke plumes and inlet flows. In [51] approximately equivariant graph neural networks are studied, and [52] study approximate equivariance in reinforcement learning. The notion of approximate equivariance was formalized by [53], who also show the importance to correctly identify the degree of equivariance (or relaxation thereof) in a given problem. In [54], a unified framework is presented to discover, enforce and promote symmetry in machine learning applications, including linear algebraic relations to find symmetry Lie-subgroups of datasets, functions and dynamic systems, given the action of a larger Lie group that fails to be symmetric.
For a review of the Kolmogorov $N$-width and recent nonlinear widths extending the concept, see the dedicated Section 2.2.
1.2 Notation
See [55] for background on differential geometry and [26] for background on Lie groups. Calligraphic letters $\mathcal{M},\,\mathcal{N}$ denote smooth manifolds. Given a point $x\in\mathcal{M}$, we let $T_{x}\mathcal{M}$ denote the tangent space at $x$ and the tangent bundle is the disjoint union $T\mathcal{M}=\dot{\bigcup}_{x\in\mathcal{M}}T_{x}\mathcal{M}$. Then $\Gamma(T\mathcal{M})$ is the set of sections of $T\mathcal{M}$ which consists of vector fields $X,Y\in\Gamma(T\mathcal{M})$ over $\mathcal{M}$. The vector field $[X,Y]$ is called the Lie-bracket of vector fields $X,Y$.
Denote as $C^{k}(\mathcal{M},\mathcal{N})$ the set of $k$-times differentiable maps from $\mathcal{M}$ to $\mathcal{N}$. We define $C^{k}(\mathcal{M}):=C^{k}(\mathcal{M},\mathbb{R})$. For $\phi\in C^{k}(\mathcal{M},\mathcal{N})$ with $k\geq 1$ the push-forward is $\phi_{*}:T_{x}\mathcal{M}\rightarrow T_{\phi(x)}\mathcal{N}$.
The group of diffeomorphisms is $\mathrm{Diff}(\mathcal{M})\subseteq C^{\infty}(\mathcal{M},\mathcal{M})$, and contains the smooth maps whose inverse is also smooth.
Further $G$ denotes a Lie group, $g,h$ denote arbitrary elements of $G$ and $e$ denotes the group identity. We denote by $\mathfrak{g}=T_{e}G$ the Lie algebra of $G$, by $\widetilde{A},\widetilde{B}\in\mathfrak{g}$ its elements, and by $\widetilde{A},\widetilde{B}$] their (left) Lie-bracket. The exponential map is $\exp:\mathfrak{g}\rightarrow G$, and we similarly denote $\exp(\widetilde{A})=:e^{\widetilde{A}}$. Lastly, $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ denotes a group action of $G$ on $\mathcal{M}$. For $x\in\mathcal{M}$ and when the choice of $\Phi$ is clear from context, we denote $\Phi_{g}(x)=g\cdot x$.
2 Background
2.1 Lie groups and their actions
Given $G$ an $n$-dimensional Lie group, and $g,h\in G$. The left and right translations are, respectively:
$L_{g}(h):=gh\,$, (4)
$R_{g}(h):=hg\,.$ (5)
Definition 2.1 (Group action)
Let $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ be a smooth map such that $\Phi_{g}(\cdot):=\Phi(g,\cdot)$ is a diffeomorphism:
$\Phi:$ $\,G\rightarrow\mathrm{Diff}(\mathcal{M})\,;\;g\mapsto\Phi_{g}(\cdot)\,.$ (6)
Then $\Phi$ is a left group action of $G$ on $\mathcal{M}$ if it is a homomorphism [26]:
$\Phi_{gh}=\Phi_{g}\circ\Phi_{h}\,$, (7)
and $\Phi$ is a right group action of $G$ on $\mathcal{M}$ if it is an anti-homomorphism [26]:
$\Phi_{gh}=\Phi_{h}\circ\Phi_{g}\,.$ (8)
An example for a left action of $G=GL(3,\mathbb{R}):=\{g\in\mathbb{R}^{3\times 3}\;|\;\det(g)\neq 0\}$ on $x\in\mathbb{R}^{3}$ is the matrix-vector product $\Phi(g,x)=gx$, and an example for a right action is $\Phi(g,x)=g^{-1}x$. Both actions represent combined rotation, scaling and sheering of $\mathbb{R}^{3}$.
The action can have a number of additional properties:
Definition 2.2 (Properties of an action)
The group action $\Phi$ can also have the following properties [26]:
  • Faithful: for all $g\in G\backslash e$ there exists $x\in\mathcal{M}$ such that $\Phi(g,x)\neq x$.
  • Free: $g\cdot x=x$ if and only if $g=e$.
  • Proper: when sequences $\{x_{n}\}$ and $\{g_{n}\cdot x_{n}\}$ converge in $\mathcal{M}$, then $\{g_{n}\}$ converges in $G$.
  • Transitive: for all $x,y\in\mathcal{M}$ there exists $g\in G$ such that $x=g\cdot y$.
Going back to the example of $GL(3,\mathbb{R})$, the left and right action are neither transitive nor free: the element $x=0\in\mathbb{R}^{3}$ can not be reached from any other $x$, and $g\cdot 0=0$ for any $g\in GL(3,\mathbb{R})$. Instead, the actions of $GL(3,\mathbb{R})$ on $\mathbb{R}^{3}\backslash 0$, i.e., $\mathbb{R}^{3}$ excluding 0, are transitive, but again not free.
Definition 2.3 (Orbit of an action)
The orbit of a point $x\in\mathcal{M}$ under the action $\Phi$ is the set
$\mathcal{O}(x):=\{g\cdot x\;|\;g\in G\}\,.$ (9)
In other words, the orbit $\mathcal{O}(x)$ collects all elements that can be reached from $x$ by applying a group element $g$. We note that $\mathcal{O}(x)$ is a submanifold of $\mathcal{M}$. Again going back to the example of $GL(3,\mathbb{R})$, $\mathcal{O}(0)=\{0\}$ for $0\in\mathbb{R}^{3}$.
Theorem 2.4 (Properties of the orbit, [26])
If $\Phi$ is transitive, then $\mathcal{O}(x)=\mathcal{M}$, and it is identical for all $x\in\mathcal{M}$. If $\Phi$ is free, then $\mathcal{O}(x)$ is isomorphic to $G$ as a manifold, for all $x\in\mathcal{M}$. If $\Phi$ is free and proper, then $\mathcal{M}/G:=\{\mathcal{O}(x)\;|\;x\in\mathcal{M}\}$ is a uniquely determined smooth manifold of dimension $\dim\mathcal{M}-\dim G$.
Definition 2.5 (Infinitesimal generator)
Given $\widetilde{A}\in\mathfrak{g}$ and $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$, the infinitesimal generator $X_{\widetilde{A}}\in\Gamma(T\mathcal{M})$ of $\widetilde{A}$ w.r.t. $\Phi$ is defined pointwise as:
$X_{\widetilde{A}}(x):=\frac{\text{d}}{\text{d}t}\Phi(\exp(\widetilde{A}t),x)\,.$ (10)
Intuitively, the infinitesimal generator $X_{\widetilde{A}}(x)$ describes the velocity at $x$ generated by the action of $\exp(\widetilde{A}t)$.
Definition 2.6 (Distribution)
A smooth distribution $\Delta\subseteq T\mathcal{M}$ assigns to any $x\in\mathcal{M}$ a subspace $\Delta_{x}\subseteq T_{x}\mathcal{M}$ smoothly, i.e., for any $x\in\mathcal{M}$ there are smooth vector fields $\{X_{1},\cdots,X_{k}\}\subseteq\Gamma(T\mathcal{M})$ and $\mathcal{U}_{x}\subseteq\mathcal{M}$ a neighborhood of $x$ such that
$\forall y\in\mathcal{U}_{x}:\;\Delta_{y}=\textnormal{span}\{X_{1}(y),\cdots,X_{k}(y)\}\,.$ (11)
A distribution is called
  • Regular of dimension $k$ if $\Delta_{x}\subseteq T_{x}\mathcal{M}$ is $k$-dimensional for every $x\in\mathcal{M}$.
  • Integrable if for any $x\in\mathcal{M}$ exists a submanifold $\mathcal{N}\subseteq\mathcal{M}$ such that $\Delta_{x}=T_{x}\mathcal{N}$.
In the present work, the relevant distribution will be the span of infinitesimal generators.
Definition 2.7 (Distribution induced by $G,\Phi$)
A Lie group $G$ and an action $\Phi$ induce, at each $x\in\mathcal{M}$, a subspace $\Delta_{x}\subseteq T_{x}\mathcal{M}$ by the image of the infinitesimal generator:
$\Delta_{x}$ $=\{X_{\widetilde{A}}(x)\,|\,\widetilde{A}\in\mathfrak{g}\}\,.$ (12)
This in turn defines a distribution $\Delta\subseteq T\mathcal{M}$ as
$\Delta=\dot{\bigcup}_{x\in\mathcal{M}}\Delta_{x}\,$, (13)
which we call the distribution induced by $\Phi$.
The induced distribution collects the tangent vectors to $\mathcal{O}(x)\subseteq\mathcal{M}$.
Theorem 2.8 (Properties of the induced distribution)
If the action $\Phi$ is free, then the induced distribution $\Delta$ is regular of dimension $\dim(G)$. The induced distribution $\Delta$ is always integrable, and tangent to the orbits $\mathcal{O}(x)$ at every $x\in\mathcal{M}$.
Proof 2.8.1
The first statement is a direct consequence of [55, Proposition 7.26]. Integrability of $\Delta$ follows from [55, Proposition 19.2] and Frobenius’ Theorem [55, Theorem 19.12]. Finally, tangency of $\Delta_{x}$ and $\mathcal{O}(x)$ is a direct consequence of Defs. 2.3, 2.5 and 2.7.
In other words, integrability of $\Delta$ directly corresponds to $\mathcal{O}(x)$ being submanifolds of $\mathcal{M}$.
2.2 The Kolmogorov N-width
Given a Hilbert space $\mathcal{M}$, and denoting by $W\subseteq\mathcal{M}$ an $N$-dimensional subspace, the Kolmogorov $N$-width of a subset $S\subseteq\mathcal{M}$ is
$d_{N}(S)=$ $\,\inf_{\begin{subarray}{c}W\subseteq\mathcal{M}\\ \dim W=N\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in W}\|x-y\|$ (14)
$=$ $\,\inf_{\begin{subarray}{c}W\subseteq\mathcal{M}\\ \dim W=N\end{subarray}}E(W,S)\,$,
with $E(W,S)=\sup_{x\in S}\;\inf_{y\in W}\|x-y\|$ the maximum approximation error.
In the context of MOR the Kolmogorov $N$-width is analyzed for the set of solution snapshots collected in $S$, and it provides a lower bound on the worst-case error of linear-subspace methods in approximating $S$. Analytic estimates of the Kolmogorov $N$-width are well-known for certain problem classes: for example, the Kolmogorov $N$-width is known to decay exponentially for diffusion problems [56], but only slowly (order $\sim\frac{1}{\sqrt{N}}$) for certain linear transport and wave problems [57]. While achieving the theoretical best case given by $d_{N}(S)$ is difficult, it can often be approached. Yet, the slow decay for linear transport suggests that linear-subspace methods are ill-suited for most transport dominated problems [15].
Lower bounds for nonlinear MOR methods are given by nonlinear alternatives to the linear-subspace Kolmogorov $N$-width. For these nonlinear widths the linear width (14) generally presents an upper bound, refer also to [58, 59].
Denoting $D_{N}:\mathbb{R}^{N}\rightarrow\mathcal{M}$ a continuous, Lipschitz decoder function, constant $\gamma>0$, then the Lipschitz $N$-width [60] of $S\subset\mathcal{M}$ is
$d^{\textnormal{Lip}}_{N}(S)=$ $\inf_{\begin{subarray}{c}D_{N},\|\cdot\|\\ \gamma\textnormal{-Lipschitz}\end{subarray}}\sup_{x\in S}\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (15)
where the final infimum is also over a suitably restricted class of norms $\|\cdot\|$ on $\mathcal{M}$.
In [61] the Sectional Komogorov $N$-width is presented. Denote by $\sigma:\mathcal{P}\rightarrow\mathcal{M}$, $\sigma\in\Gamma$ a section of a fiber bundle with the parameter manifold $\mathcal{P}$ as the base space, and the state manifold $\mathcal{M}$ as the fiber. Then the linear Sectional $N$-width measures distance to an optimally chosen linear sub-bundle $\Gamma_{N}$:
$d^{\textnormal{Sec}}_{N}(S)=\inf_{\begin{subarray}{c}\Gamma_{N}\subset\Gamma\;\\ \dim\Gamma_{N}\leq N\\ \textnormal{linear}\end{subarray}}\sup_{\mu\in\mathcal{P}}\inf_{\sigma_{N}\in\Gamma_{N}}\|\sigma(\mu)-\sigma_{N}(\mu)\|\,.$ (16)
Non-linear extensions in the same work [61] emphasize that $\mathcal{M}$ can be a manifold, in which case $\Gamma_{N}$ is a sub-bundle for which the fiber $\sigma_{N}(\mu)$ is a submanifold of $\sigma(\mu)$.
3 Problem formulation
Assume the FOM to be the dynamic system on $\mathcal{M}$:
$\dot{x}=X_{\mu}(x)\,,\;x(0)=x_{\mu,0}\,$, (17)
with parameters $\mu\in\mathcal{P}$ parameterizing both the vector field $X_{\mu}\in\Gamma(T\mathcal{M})$ and initial conditions $x_{\mu,0}$. Solutions of (17) will be denoted $x(t)$, or $x_{\mu}(t)$ to emphasize the parameter-dependence. Let a set of solution snapshots be given as
$S=\{x_{\mu}(t)\,|\,t\in[0,T],\mu\in\mathcal{P}\}\,.$ (18)
We will mention it explicitly when finite or countably infinite snapshot sets $S$ are considered, or when $S\subset\mathcal{M}\times\mathcal{P}\times[0,T]$ by abuse of notation.
The goal of this paper is to find a Lie group $G$ with dimension $\dim G\ll\dim\mathcal{M}$, action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$, and dynamics of $g_{\mu}(t)$ such that
$x_{\mu}(t)\approx\Phi(g_{\mu}(t),x_{\mu,0})$ (19)
for $x_{\mu}(t)\in S$. In particular, we aim to answer the questions:
  1. 1.
    How can a dynamics for $g_{\mu}(t)$ be chosen and optimized? In Section 4.1.
  2. 2.
    What is a theoretical error bound for the resulting methods? In Section 4.2.
  3. 3.
    What are choices for $G$, $\Phi$ such that dynamics on $G$ can be evaluated more efficiently than the FOM dynamics? In Section 5.
  4. 4.
    How does this approach relate to submanifold methods? In Appendix D.
4 MOR via Lie groups
We describe MORLie and introduce the group Kolmogorov $N$-width.
4.1 ROM dynamics
We investigate the choice of dynamics for $g_{\mu}(t)$ in (19) to arrive at a description of the ROM dynamics in MORLie.
We want to choose the Lie group dynamics such that the evolution of $\Phi(g_{\mu}(t),x_{\mu,0})$ closely follows (w.r.t. a to-be-defined metric) the full-order state $x_{\mu}(t)$. Since the FOM dynamics $X_{\mu}(x_{\mu})$ vary with $x_{\mu}$, the optimal Lie group dynamics should also vary with $x_{\mu}$. We encode this by a map $\rho_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$, and immediately state the main technical result of this subsection: the induced dynamics on $\mathcal{M}$ generated by dynamics on $G$. The key idea is that a vector in the Lie algebra produces a vector field on $\mathcal{M}$ via the infinitesimal generator.
Theorem 4.1 (MorLie reduction, reconstruction and induced dynamics)
Given a Lie group $G$, a left action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ and a map $\rho_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$. Define $\bar{x}_{\mu}(t)\in\mathcal{M}$ by
$\bar{x}_{\mu}(t):=\Phi(g_{\mu}(t),x_{\mu,0})\,$, (20)
and define dynamics on $G$ to follow $\rho_{\mu}(\bar{x}_{\mu})\in\mathfrak{g}$:
$\dot{g}_{\mu}={R_{g_{\mu}}}_{*}\rho_{\mu}\big(\bar{x}_{\mu}\big)\,,\;g(0)=e\,.$ (21)
Then
$\frac{\text{d}}{\text{d}t}\bar{x}_{\mu}(t)=X_{\rho_{\mu}\big(\bar{x}_{\mu}(t)\big)}\big(\bar{x}_{\mu}(t)\big)\,$, (22)
where $X_{\rho_{\mu}(\bar{x}_{\mu})}\in\Gamma(\Delta)$ is the infinitesimal generator of $\rho_{\mu}(\bar{x}_{\mu})$ w.r.t. $\Phi$ (cf. Def. 2.5), and $\Delta\subseteq T\mathcal{M}$ is the induced distribution (cf. Def. 2.7). Equation (22) also holds when $\Phi$ is a right action, if the left-translation were used in (21).
Proof 4.1.1
For a left action $\Phi$, differentiation of $\bar{x}_{\mu}(t)$ yields:
$\frac{\text{d}}{\text{d}t}\bar{x}_{\mu}(t)$ $=\frac{\text{d}}{\text{d}t}\Phi(g_{\mu}(t),x_{\mu,0})$ (23)
$=\frac{\text{d}}{\text{d}s}\Phi(e^{\rho(\bar{x}_{\mu}(t))s}g_{\mu}(t),x_{\mu,0})$ (24)
$=\frac{\text{d}}{\text{d}s}\Phi(e^{\rho(\bar{x}_{\mu}(t))s},\Phi_{g_{\mu}(t)}x_{\mu,0})$ (25)
$=X_{\rho_{\mu}(\bar{x}_{\mu}(t))}\big(\bar{x}_{\mu}(t)\big)$ (26)
The second equality uses that $g_{\mu}(s)$ in (21) is tangent to $e^{\rho(\bar{x}_{\mu}(t))s}g_{\mu}(t)$ at $t=s$. The third equality uses (7), using that $\Phi$ is a left action. The fourth equality uses (10). For a right action $\Phi$, and if $R_{g_{\mu}}$$_{*}$ in (21) were replaced by $L_{g_{\mu}}$$_{*}$: then $g_{\mu}(s)$ would be tangent to $g_{\mu}(t)e^{\rho(\bar{x}_{\mu}(t))s}$ at $t=s$, in the second equality, and the third equality would use (8).
We formally define further terms inspired by Theorem 4.1:
Definition 4.2 (MorLie reduction & approximated dynamics)
Given the FOM $(\mathcal{M},X_{\mu})$, and a tuple $(G,\Phi,\rho_{\mu})$ of a Lie group $G$, a group action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ and $\rho_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$. Then we call $\rho_{\mu}$ the reduced vector field, and $\bar{X}\in\Gamma(\Delta)\subseteq\Gamma(T\mathcal{M})$
$\bar{X}_{\mu}(x):=X_{\rho_{\mu}(x)}(x)\,$, (27)
the approximated dynamics. This defines the dynamical system
$\dot{\bar{x}}(t)=\bar{X}_{\mu}(\bar{x})\,,\;\bar{x}(0)=x_{\mu,0}\,.$ (28)
We call $\bar{x}(t)$ (cf. (20)) the reconstructed solution.
Thus, $\rho$ defines dynamics on $G$, which induce approximated dynamics $\bar{X}$ on $\mathcal{M}$, constrained to the orbits $\mathcal{O}(x)$ and tangent to the distribution $\Delta$.
A reduced order model is obtained when $\dim G<\dim\mathcal{M}$: then Theorem (4.1) lets us solve a low dimensional dynamical system on the Lie group $G$ to compute solutions $\bar{x}_{\mu}$ of the high-dimensional approximated dynamics $\bar{X}$ on $\mathcal{M}$. We call (21) a family of reduced order models on the Lie group $G$, since they vary with initial state $x_{\mu,0}$ and parameter $\mu$. This presents a formulation for MOR on manifolds that allows to approximate solutions globally, where different $x_{\mu,0}$ can lead to different dynamical systems on $G$. Distinct ROMs can in principle also be qualitatively different, i.e., having different stability properties, number of equilibria or periodic orbits.
Theorem 4.1 shows that the evolution of $\bar{x}_{\mu}(t)$ in (21) follows an approximated dynamics $\bar{X}_{\mu}\in\Gamma(\Delta)$, which is restricted to lie within the distribution $\Delta\subseteq T\mathcal{M}$ induced by $G,\Phi$ (see Definition 2.7). Theorem 4.1 can alternatively be interpreted as a choice of a family of approximate dynamics $\bar{X}_{\mu}\in\Gamma(\Delta)$ that admit a family of ROMs of the form (21).
We emphasize that (21) represents any autonomous first-order dynamics $\dot{g}_{\mu}=f(g_{\mu})$ without loss of generality: the choice of dynamics $\rho_{\mu}(\bar{x}_{\mu})\in\mathfrak{g}$ can be written as $\widetilde{A}_{\mu}(g_{\mu}):=\rho_{\mu}(\bar{x}_{\mu})\in\mathfrak{g}$, hiding dependence on $x_{\mu,0}$ on the RHS in dependence on $\mu$ on the LHS. Then $\widetilde{A}_{\mu}(g_{\mu})={R_{g^{-1}}}_{*}f(g_{\mu})$ implements arbitrary autonomous first-order dynamics.
Theorem 4.1 also shows that the approximated solution $\bar{x}(t)\in\mathcal{O}(x_{0})$ is restricted to lie in the orbit $\mathcal{O}(x_{0})\subseteq\mathcal{M}\,$ (see Definition 2.3). The following Section uses this fact to define the group Kolmogorov $N$-width.
4.2 The group Kolmogorov N-width
Suppose that $\mathcal{M}$ is a metric space equipped with a distance $\textnormal{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}$, and given the set $S\subseteq\mathcal{M}$ representing solution snapshots (cf. (18)). Define $\Xi_{N}$ as a set of pairs $(G,\Phi)\in\Xi_{N}$ of $N$-dimensional Lie groups $G$ and actions $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$, henceforth called admissible pairs. Then we define:
Definition 4.3 (Kolmogorov $\Xi_{N}$-width)
The Kolmogorov $\Xi_{N}$-width of a set $S\subseteq\mathcal{M}$ relative to an arbitrary point $x_{0}\in S$ is
$d_{\Xi_{N}}(S,x_{0}):=\inf_{\begin{subarray}{c}(G,\Phi)\in\Xi_{N}\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in\mathcal{O}(x_{0})}\textnormal{dist}(x,y)\,.$ (29)
For $S$ a set of solution snapshots (18), and denoting $x_{\mu}([a,b]):=\{x_{\mu}(t)\;|\;t\in[a,b]\}$, we define the Kolmogorov $\Xi_{N}$-width of $S$ over the initial conditions $x_{0,\mu}\in S$ as
$d^{T}_{\Xi_{N}}(S):=\sup_{\mu\in\mathcal{P}}d_{\Xi_{N}}\big(x_{\mu}([0,T]),x_{0,\mu}\big)\,$, (30)
and the Kolmogorov $\Xi_{N}$-width of $S$ over a time-horizon $\tau\leq T$ as
$d^{T,\tau}_{\Xi_{N}}(S):=\sup_{\begin{subarray}{c}\mu\in\mathcal{P},\\ t\in[0,T-\tau]\end{subarray}}d_{\Xi_{N}}\big(x_{\mu}([t,t+\tau]),x_{\mu}(t)\big)\,.$ (31)
Intuitively, the group width measures how well solution trajectories can be approximated by a single group orbit through the initial state.
In particular, $d_{\Xi_{N}}(S,x_{0})$ is the worst case error over the orbit $\mathcal{O}(x_{0})$, for the best choice of admissible pair $(G,\Phi)\in\Xi_{N}$. In turn, $d^{T}_{\Xi_{N}}(S)$ formalizes the error between FOM-trajectories $x_{\mu}(t)$ and ROM-trajectories $\bar{x}_{\mu}(t)$ (cf. (19)), returning the worst case over all initial conditions. Finally, $d^{T,\tau}_{\Xi_{N}}(S)$ measures a worst-case error between $x_{\mu}(t)$ and $\bar{x}_{\mu}(t)$ over a time-horizon $\tau$, and without restricting itself to ROM-trajectories starting at $x_{\mu,0}$. Alternatively, we can interpret $\frac{1}{\tau}d^{T,\tau}_{\Xi_{N}}(S)$ as a bound on the increase of $\textnormal{dist}\big(\bar{x}(t),x(t)\big)$ per $\tau$. We note that Definitions (30) and (31) coincide for $\tau=T$ ($d^{T}_{\Xi_{N}}(S)=d^{T,T}_{\Xi_{N}}(S)$), in which case the supremum over $t\in[0,T-\tau]$ in (31) is over a single point $t=0$.
If $\Xi_{N}$ is not restricted, e.g., $\Xi_{1}$ contains all possible pairs of one-dimensional Lie groups and compatible actions, then $d^{T,\tau}_{\Xi_{1}}(S)=0$ irrespective of $S$: this can be seen for $G=(\mathbb{R}^{1},+)$ and $\Phi(a,x)=\Psi^{a}_{X}(x)$ with $\Psi^{t}_{X}:\mathcal{M}\rightarrow\mathcal{M}$ the flow of $X\in\Gamma(TM)$. This exactly recovers the FOM trajectories, but does not result in a useful ROM, as this action is computationally intense to evaluate.
Similarly, if $y$ were not restricted to $y\in\mathcal{O}(x_{0})$ then $d_{\Xi_{N}}(S,S)=0$, i.e., it would be possible to pick $y=x$ for any $x\in S$.
Proposition 4.4
Assume that $\mathcal{M}$ is a Hilbert space with $\textnormal{dist}(x,y)=\|x-y\|$, define
$\Xi^{\textnormal{Vec}}_{N}:=\{$ $(\mathbb{R}^{N},\Phi)\;|\;\Phi(g,x):=x+g^{i}e_{i}\,$, (32)
$\{e_{1},\cdots,e_{N}\}\subseteq\mathcal{M}\textnormal{ lin. indep.}\}\,.$
Then we recover the Kolmogorov $N$-width (14) from (29)
$d_{N}(S)=d_{\Xi^{\textnormal{Vec}}_{N}}(S,0)\,.$ (33)
Proof 4.4.1
Different orbits $\mathcal{O}(x_{0})$ (cf. Def. 2.3) generated by $(\mathbb{R}^{N},\Phi)\in\Xi^{\textnormal{Vec}}_{N}$ correspond to $N$-dimensional affine subspaces through $x_{0}$:
$\mathcal{O}_{(\mathbb{R}^{N},\Phi)}(x_{0})=\{x_{0}+g^{i}e_{i}\;|\;g\in\mathbb{R}^{N}\}\,.$ (34)
Thus, the orbits $W=\mathcal{O}_{(\mathbb{R}^{N},\Phi)}(0)$ correspond to $N$-dimensional subspaces. Further, given any $N$-dimensional subspace $W$ spanned by $\{e_{1},\cdots,e_{N}\}\subseteq W$, there is $(\mathbb{R}^{N},\Phi)\in\Xi^{\textnormal{Vec}}_{N}$ such that $W=\mathcal{O}_{(\mathbb{R}^{N},\Phi)}(0)$. Hence, the first infimum in $d_{\Xi^{\textnormal{Vec}}_{N}}(S,0)$ is equivalent to an infimum over subspaces $W\subseteq\mathcal{M}$ such that $\dim W=N$, and the second infimum is over elements $v\in W=\mathcal{O}(0)$:
$d_{\Xi^{\textnormal{Vec}}_{N}}(S,0)$ $=\inf_{\begin{subarray}{c}(G,\Phi)\in\Xi^{\textnormal{Vec}}_{N}\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in\mathcal{O}(0)}\textnormal{dist}(x,y)$ (35)
$=\inf_{\begin{subarray}{c}W\subseteq\mathcal{M}\\ \dim W=N\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in W}\|x-y\|\,.$ (36)
For $\mathcal{M}$ a vector space, including linear, affine or even nonlinear transformations is possible by considering different choices of $\Xi_{N}$, such as
$\Xi^{\textnormal{Lin}}_{N}$ $:=\{(G,\Phi)\;|\;\dim G=N\,,\;\Phi\textnormal{ linear}\}\,$, (37)
$\Xi^{\textnormal{Aff}}_{N}$ $:=\{(G,\Phi)\;|\;\dim G=N\,,\;\Phi\textnormal{ affine}\}\,$, (38)
$\Xi^{\textnormal{Pr}}_{N}$ $:=\{(G,\Phi)\;|\;\dim G=N\,,\;\Phi\textnormal{ proper}\}\,.$ (39)
In particular, linear $\Phi$ are linear representations and affine $\Phi$ are affine representations such that the $N$-dimensional Lie groups in $\Xi^{\textnormal{Lin}}_{N}$ and $\Xi^{\textnormal{Aff}}_{N}$ can always be associated with $N$-dimensional subsets of general linear group $GL(n,\mathbb{R})$, general affine group $\textnormal{Aff}(n,\mathbb{R})$ respectively, with $n=\dim\mathcal{M}$. It can also be shown that $\Xi^{\textnormal{Pr}}_{N}$ includes $\Xi^{\textnormal{Vec}}_{N}\,,\,\Xi^{\textnormal{Lin}}_{N}\,,\,\Xi^{\textnormal{Aff}}_{N}$ as subsets, so it follows that
$d_{\Xi^{\textnormal{Pr}}_{N}}(S,0)\leq d_{N}(S)\,.$ (40)
Therefore, group methods can break the Kolmogorov barrier. The set $\Xi^{\textnormal{Aff}}_{N}$ includes $\Xi^{\textnormal{Lin}}_{N}$ and $\Xi^{\textnormal{Vec}}_{N}$ as (non-intersecting) subsets, so it also follows that
$d_{\Xi^{\textnormal{Aff}}_{N}}(S,0)\leq d_{\Xi^{\textnormal{Lin}}_{N}}(S,0)\,$, (41)
$d_{\Xi^{\textnormal{Aff}}_{N}}(S,0)\leq d_{\Xi^{\textnormal{Vec}}_{N}}(S,0)\,.$ (42)
Yet, affine MOR only provides a marginal advantage over linear MOR since $\Xi^{\textnormal{Lin}}_{N+1}$ includes $\textnormal{Aff}_{N}$, such that $d_{\Xi^{\textnormal{Lin}}_{N+1}}(S,0)\leq d_{\Xi^{\textnormal{Aff}}_{N}}(S,0)$. Sections 6, 7 will show more significant examples of breaking the Kolmogorov barrier by means of group methods, and Section 6 provides a specific group Kolmogorov $N$-width of linear transport.
Remark 4.5
It is interesting to restrict $\Xi_{N}$ to pairs $(G,\Phi)$ such that the $\Phi$ is of a low computational complexity. Possible choices are the aforementioned linear and affine transformations, but also nonlinear actions or flows of differential equations could be parameterized by shallow neural networks to provide interesting classes.
Remark 4.6
When the action of a pair $(G,\Phi)\in\Xi_{N}$ is transitive, i.e., $\mathcal{O}(x_{0})=\mathcal{M}$, we likewise get that $d_{\Xi_{N}}(S,x_{0})=0$. However, for proper and transitive actions (ignoring improper transitive actions resulting from topologically transitive flows) we have $\dim G>\dim\mathcal{M}$, which makes the case uninteresting for MOR, but leads to the interesting class of Lie group methods for integration of lifted (rather than reduced) dynamics on $G$.
The Lipschitz $N$-width (15) can also be recovered from (29):
Proposition 4.7
Assume that $\mathcal{M}$ is a Hilbert space with $\textnormal{dist}(x,y)=\|x-y\|$, $T^{N}=S^{1}\times\cdots\times S^{1}$ is the $N$-torus. Let $D_{N}:\mathbb{R}^{N}\rightarrow\mathcal{M}$ denote the decoder function in (15), with constant $\gamma>0$ and $\|\cdot\|$ from a suitably restricted class of norms on $\mathcal{M}$. Also construct a discontinuous embedding $\textnormal{Emb}:T^{N}\rightarrow\{x\in\mathbb{R}^{N}\;|\;\|x\|\leq B\}$ mapping $N$ copies of $S^{1}$ to $N$ copies of the half-open interval $(-B,B]\in\mathbb{R}$. Define
$\Xi^{\textnormal{Lip}}_{N}:=\{$ $(T^{N},\Phi)\;|\;\Phi(g,x):=x+D_{N}(\textnormal{Emb}(g))\,$, (43)
$D_{N}\textnormal{is }\gamma\textnormal{-Lipschitz}\}\,.$
Then we recover the Lipschitz $N$-width (15) from (29)
$d^{\textnormal{Lip}}_{N}(S)=\inf_{\|\cdot\|_{\mathcal{M}}}d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)\,.$ (44)
Proof 4.7.1
The orbit $\mathcal{O}(0)$ is
$\mathcal{O}_{(T^{N},\Phi)}(0)=D_{N}(\textnormal{Emb}(T^{N}))\subset\mathcal{M}\,.$ (45)
Hence, the first infimum in $d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)$ is
$\inf_{y\in\mathcal{O}(0)}\|x-y\|$ $=\inf_{g\in T^{N}}\|x-D_{N}(\textnormal{Emb}(g))\|$ (46)
$=\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (47)
where the final step holds since the image of Emb is dense in the closed ball $\{x\;|\;\|x\|\leq B\}$, and we recover
$\inf_{\|\cdot\|}d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)$ $=\inf_{\begin{subarray}{c}(G,\Phi)\in\Xi^{\textnormal{Lip}}_{N}\\ \|\cdot\|\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in\mathcal{O}(0)}\textnormal{dist}(x,y)$ (48)
$=\inf_{\begin{subarray}{c}D_{N},\|\cdot\|\\ \gamma\textnormal{-Lipschitz}\end{subarray}}\;\sup_{x\in S}\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (49)
which is equivalent to (15).
Here, the Lie group formalism is taken very loosely: the map $\Phi(x,g)=x+D_{N}(\textnormal{Emb}(g))$ in (43) is not explicitly restricted to be a homomorphism (cf. Def. 2.1), with the result that properties of the orbits (cf. Def. 2.4) and induced distribution (cf. Def. 2.8) are no longer guaranteed. The resulting orbits may self-intersect, thus overparameterizing $\mathcal{M}$ – this problem would be avoided if the homomorphism property can be enforced.
A ROM in MORLie can also be evaluated by means of the Sectional $N$-width (16).
Proposition 4.8
Assume that $\mathcal{M}$ is a Hilbert space with $\textnormal{dist}(x,y)=\|x-y\|$. Let $\Phi_{\mu}(g,x)$ be a parameter-dependent action. Define
$\Xi^{\textnormal{Sec}}_{N}:=\{$ $(\mathbb{R}^{N},\Phi_{\mu})\;|\;\Phi_{\mu}(g,x):=x+g^{i}e_{\mu,i}\,$, (50)
$\{e_{\mu,1},\cdots,e_{\mu,N}\}\subseteq\mathcal{M}\textnormal{ lin. indep.}\}\,.$
Any such action $\Phi_{\mu}$ induces a linear section
$\sigma_{N}(\mu)=\mathcal{O}_{(\mathbb{R}^{N},\Phi_{\mu})}(x_{\mu})\,.$ (51)
$d^{\textnormal{Sec}}_{N}(S)=d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)\,.$ (52)
I.e., a parameterized linear action gives rise to a linear section $\sigma_{N}(\mu)$ with varying linear fiber. Similarly, nonlinear sections can be induced by general parameterized actions.
5 Optimization in MORLie
We want to choose $(G,\Phi,\rho_{\mu})$ to arrive at ROM dynamics that can be evaluated more efficiently than the FOM dynamics, and enable reconstructed solutions to approximate FOM solutions (cf. (19)). To this end, intrusive and non-intrusive methods for MORLie are presented, also preparing for the examples in Sections 6, 7. In both cases we cast the process of determining $(G,\Phi,\rho_{\mu})$ for a given $(\mathcal{M},X_{\mu},S)$ into the form of an optimization problem.
Let $\theta\in\mathbb{R}^{n_{\theta}}$ denote coefficients of a parameterization $\rho:\mathbb{R}^{n_{\theta}}\times\mathcal{P}\rightarrow C^{\infty}(\mathcal{M},\mathfrak{g})$, i.e., $\rho_{\theta,\mu}:\mathcal{M}\rightarrow\mathfrak{g}$ describes a subset $\Upsilon\subseteq C^{\infty}(\mathcal{M},\mathfrak{g})$ of reduced vector fields. Similar to Section 4.2, let $\Xi$ be a to-be-determined set of pairs $(G,\Phi)$, called admissible pairs, and let $J$ be a to-be-determined real-valued cost-function. The general form of the optimization problem will be:
$(G^{*},\Phi^{*},\theta^{*})=\arg\min_{\begin{subarray}{c}(G,\Phi)\in\Xi\,,\\ \theta\in\mathbb{R}^{n_{\theta}}\end{subarray}}J\,.$ (53)
In the following subsections we elaborate on (53), treating in more detail choices of $\Xi$ and $\Upsilon$, a choice of $J$ for intrusive and non-intrusive optimization of the ROM. Finally, we will present a strategy for solving (53) for $(G^{*},\Phi^{*},\theta^{*})$, and further dimensionanility reduction going from $(G^{*},\Phi^{*},\rho_{\theta,\mu}^{*})$ to $(H^{*},\Phi^{*},\rho_{\theta,\mu}^{*})$ with $H^{*}\subset G^{*}$.
5.1 Search spaces for $(G,\Phi)$
We show how a search space $\Xi$ for $(G,\Phi)$ may be determined, for a given manifold $\mathcal{M}$.
A full classification of Lie group actions with particular properties is an open research issue [62], with non-trivial dependence on the manifold $\mathcal{M}$. For example, not every Lie group $G$ has a meaningful action on a given manifold $\mathcal{M}$: there is a free, faithful and proper action of $SO(3)$ on $\mathbb{R}^{3}$ but not on $T^{3}$. There is a free action of $SU(2)$ on $S^{7}$, but not on $S^{5}$. Also the types of possible actions $\Phi$ depend on $\mathcal{M}$: for example, isometric actions are only defined on Riemannian manifolds, and symplectic actions are only defined on symplectic manifolds. Linear and affine actions can only be defined when $\mathcal{M}$ is a vector-space, in which case actions are referred to as representations, leading to a rich and ongoing field of representation theory [63]. We make no attempt to summarize the vast field of possible pairings of manifolds, Lie groups and actions, and relegate the interested reader to [62, 63, 64].
The possible search spaces $\Xi$ for $(G,\Phi)$ non-trivially depend on $\mathcal{M}$, since both $G,\Phi$ non-trivially depend on $\mathcal{M}$.
In the following, we make minor remarks on possible choices of $\Xi$, assuming prior information of a Lie group $G$ and action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ that act on $\mathcal{M}$. Then a set $\Xi$ of admissible Lie group and action pairs can be induced:
Definition 5.1 (Induced admissible pairs)
Given a manifold $\mathcal{M}$, Lie group $G$ and group action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$, define the set of admissible pairs induced by $(G,\Phi)$:
$\Xi^{G,\Phi}:=\{(H,\Phi)\;|\;H\subseteq G\}\,$, (54)
where $H\subseteq G$ denotes a Lie subgroup. Then induced sets of admissible pairs of a fixed dimension $N$, or with additional properties (cf. Def. 2.2) can be defined as
$\Xi^{G,\Phi}_{N}$ $:=\{(H,\Phi)\;|\;H\subseteq G\,,\dim H=N\}\,.$ (55)
$\Xi^{G,\Phi}_{\textnormal{Pr}}$ $:=\{(H,\Phi)\;|\;H\subseteq G\textnormal{ proper subgroup}\}\,.$ (56)
Similar definitions allow to preserve any properties of the action, further pruning the set of subgroups and action pairs $\Xi^{G,\Phi}$.
We consider a few special cases to apply Definition 5.1. For example, given $\mathcal{M}=\mathbb{R}^{3}$, the affine group $Aff(3)$ (see Sec. 7.2) and its action on $\mathbb{R}^{3}$ induce $\Xi^{G,\Phi}$ containing also $SE(n)$, $SO(n)$, $(\mathbb{R}^{n},+)$ (for $n\leq 3$) as proper subgroups.
For arbitrary manifolds $\mathcal{M}$, the diffeomorphism group $\mathrm{Diff}(\mathcal{M})$ of smooth invertible maps $\varphi:\mathcal{M}\rightarrow\mathcal{M}$ (a topological Lie group) induces a versatile search space $\Xi^{G,\Phi}$, whose subgroups were described in terms of neural nets parameterizing vector fields, in [29].
Homogenous manifolds $\mathcal{M}$ are a class of manifolds for which a transitive action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ is already known. Then the induced distribution $\Delta=T\mathcal{M}$ (cf. Def. 2.7) already contains the FOM dynamics $X\in\Gamma(T\mathcal{M})$, and it is guaranteed that $\Xi^{G,\Phi}$ contains a group and action pair such that the ROM exactly describes the FOM. We will reuse this fact in Section 5.5.
5.2 Search space for reduced vector fields
We consider two cases for parameterizing $\rho_{\theta,\mu}:\mathcal{M}\rightarrow\mathfrak{g}$, corresponding to the search space $\Upsilon\subseteq C^{\infty}(\mathcal{M},\mathfrak{g})$.
First, it is possible to directly parameterize the map $\rho_{\theta,\mu}:\mathcal{M}\rightarrow\mathfrak{g}$ in local charts $(U_{i},Q_{i})$ with $U_{i}\subseteq\mathcal{M}$ and $Q_{i}:U_{i}\rightarrow\mathbb{R}^{n}$. To this end, let there be a finite collection of charts $(U_{i},Q_{i})$ that cover $\mathcal{M}$ and have smooth transition maps, and given a partition of unity $\sigma_{i}:\mathcal{M}\rightarrow\mathbb{R}$ with respect to the charts, i.e., functions such that $\sigma_{i}(x)>0$ for $x\in U_{i}$, $\sigma_{i}(x)=0$ for $x\notin U_{i}$ and $\sum_{i}\sigma_{i}(x)=1$ for all $x\in\mathcal{M}$. Further, let $\Lambda:\mathbb{R}^{\dim\mathfrak{g}}\rightarrow\mathfrak{g}$ be an invertible, linear map implementing a basis of the Lie algebra. Then $\rho_{\theta,\mu}$ can be parameterized in terms of component functions $f^{i}_{\theta}:\mathbb{R}^{\dim\mathcal{M}}\times\mathbb{R}^{\dim\mathcal{P}}\rightarrow\mathbb{R}^{\dim\mathfrak{g}}$ as
$\rho_{\theta,\mu}(x)=\sum_{i}\Lambda\big((\sigma_{i}f^{i}_{\theta})(Q_{i}^{-1}(x),\mu)\big)\,.$ (57)
Component functions $f^{i}_{\theta}$ can be implemented e.g., by autoencoders, series expansions, or other parameterizations. However, direct parameterization of $\rho_{\theta,\mu}$ will require explicit computation of $\bar{x}=\Phi(g_{\mu}(t),x_{\mu,0})$ at every time-instance, possibly incuring a large computational overhead by working with the FOM state directly when the ROM dynamics are solved.
As an alternative, we parameterize $\rho_{\theta,\mu}$ such that the reduced-order state $g_{\mu}$ (rather than the full-order state $\bar{x}$) can be used to evaluate the ROM dynamics. Removing reference to the full-order state reduces computational overhead, and is typically referred to as hyperreduction. Let there be a finite collection of charts $(\widetilde{U}_{i},\widetilde{Q}_{i})$ that cover $G$ and have smooth transition maps, and given a partition of unity $\widetilde{\sigma}_{i}:G\rightarrow\mathbb{R}$ with respect to the charts. Then $\rho_{\theta,\mu}(\Phi(g_{\mu},x_{\mu,0}))$ can be parameterized in terms of component functions $f^{i}_{\theta}:\mathbb{R}^{\dim G}\times\mathbb{R}^{\dim\mathcal{P}}\rightarrow\mathbb{R}^{\dim\mathfrak{g}}$ as
$\rho_{\theta,\mu}(\Phi(g_{\mu},x_{\mu,0}))=\sum_{i}\Lambda\big((\widetilde{\sigma}_{i}\widetilde{f}^{i}_{\theta})(\widetilde{Q}_{i}^{-1}(g_{\mu}),\mu)\big)\,.$ (58)
The parameterizations are otherwise equivalent in terms of approximation-power: no information is lost by not considering $\bar{x}$ directly, since all relevant complexity w.r.t. different initial conditions $x_{\mu,0}$ is implicit in the fixed parameter $\mu$, and the full-order state is fully determined by $g_{\mu},x_{\mu,0}$. The component functions $\widetilde{f}^{i}_{\theta}$ can be implemented similarly to $f^{i}_{\theta}$. For high-dimensional parameters $\mu$ it may also be interesting to further reduce them to their most important features, e.g., by use of autoencoders.
5.3 Intrusive and non-intrusive MORLie
For a first intrusive cost-function, we consider a smooth Riemannian manifold $\mathcal{M}$ with metric-induced norm $\|\cdot\|$ on local tangent-spaces $T_{x}\mathcal{M}$, and metric-induced volume form. With a finite parameter set $\mathcal{P}_{f}\subset\mathcal{P}$ we define
$J$ $(G,\Phi,\rho_{\theta,\mu},X_{\mu})$ (59)
$=\sum_{\mu_{i}\in\mathcal{P}_{f}}\int_{\mathcal{M}}\|X_{\mu_{i}}(x)-X_{\rho_{\theta,\mu_{i}}(x)}(x)\|\,.$
In (59), it is the explicit dependence on $X_{\mu_{i}}$ that marks the cost-function as intrusive.
Theorem 5.2
For fixed $G,\Phi$ the optimization problem (53) with cost (59) has an explicit solution $\rho^{*}_{\mu}\in C^{\infty}(\mathcal{M},\mathfrak{g})$ given by
$\rho^{*}_{\mu}(x)={X_{\mathfrak{g}}}^{\dagger}\Pi_{\Delta}X_{\mu_{i}}(x)\,$, (60)
where $\Pi_{\Delta}:T\mathcal{M}\rightarrow\Delta$ is the metric-projection onto the induced distribution $\Delta$ (cf. Def. 2.7) and ${X_{\mathfrak{g}}}^{\dagger}:\Delta\rightarrow\mathfrak{g}$ inverts the infinitesimal generator ($\forall{\widetilde{A}\in\mathfrak{g}}:{X_{\mathfrak{g}}}^{\dagger}X_{\widetilde{A}}=\widetilde{A}$) [^1] The operator ${X_{\mathfrak{g}}}^{\dagger}\Pi_{\Delta}$ is a coordinate-free version of a Moore-Penrose pseudo-inverse of $X_{\mathfrak{g}}$, i.e., corresponding to algebraic inversion.. Further, if $X_{\mu}$ is smooth, then so is $\rho^{*}_{\mu}$.
Proof 5.2.1
The proposed explicit solution (60) minimizes the integrand $\|X_{\mu_{i}}(x)-X_{\rho_{\theta,\mu_{i}}(x)}(x)\|$ for all $x\in\mathcal{M},\mu_{i}\in\mathcal{P}$. Thus, also (59) is minimized. Finally, $\rho^{*}_{\mu}$ is the composition of smooth maps $X_{\mathfrak{g}}^{\dagger}$ and $\Pi_{\Delta}$, and the vector field $X_{\mu}$ which is smooth by assumption, so $\rho^{*}_{\mu}$ is also smooth.
The result of Theorem 5.2 extends to $\theta$ dependent $\rho_{\theta,\mu}$ if there are $\theta^{*}$ such that $\rho_{\theta^{*},\mu}=\rho_{\mu}^{*}$.
For a discrete implementation consider a finite set of state-parameter snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$ and parameters $\mu_{i}$:
$S:=\{\big(x_{i,k},\mu_{i}\big)\;|\;k\in\mathcal{I}_{k},i\in\mathcal{I}_{i}\}\,.$ (61)
We define
$J$ $(G,\Phi,\rho_{\theta,\mu},S,X_{\mu})$ (62)
$=\sum_{(x_{i,k},\mu_{i})\in S}\|X_{\mu_{i}}(x_{i,k})-X_{\rho_{\theta,\mu_{i}}(x_{i,k})}(x_{i,k})\|\,.$
Contrary to (59), the cost (62) can also be implemented as a non-intrusive cost-function, by replacing $X_{\mu_{i}}(x_{i,k})$ with velocity-measurements.
Theorem 5.3
For fixed $(G,\Phi)$, the optimization problem (53) with cost (62) has the non-unique solution $\rho^{*}_{\mu}\in C^{\infty}(\mathcal{M},\mathfrak{g})$ given by (60).
Proof 5.3.1
The proposed explicit solution (60) minimizes the summand $\|X_{\mu_{i}}(x_{i,k})-X_{\rho_{\theta,\mu_{i}}(x_{i,k})}(x_{i,k})\|$ for all $\mu_{i},x_{i,k}$, and thus also the sum. However, values of $\rho_{\mu}(x)$ for $(x,\mu)\notin S$ do not influence the value of the cost (62), so they are not uniquely determined by the optimization problem.
The result of Theorem 5.3 extends to $\theta$-dependent $\rho_{\theta,\mu}$ if there are $\theta^{*}$ such that $\rho_{\theta^{*},\mu}=\rho_{\mu}^{*}$. Such a parametrization is generally easier to find than for Theorem 5.2, since the equality only has to hold at a finite number of points.
For a velocity-free non-intrusive cost-function we consider a smooth space-time manifold $\mathcal{M}\times\mathbb{R}$ with spatial metric $\textnormal{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, and finite set of state-time-parameter snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$, time-instances $t_{i,k}$ and parameters $\mu_{i}$:
$S:=\{\big(x_{i,k},t_{i,k},\mu_{i}\big)\;|\;k\in\mathcal{I}_{k},i\in\mathcal{I}_{i}\}\,.$ (63)
We define
$J$ $(G,\Phi,\rho_{\theta,\mu},S)$ (64)
$=\sum_{(x_{i,k},t_{i,k},\mu_{i})\in S}\textnormal{dist}\big(x_{i,k+1},\Phi(e^{\rho_{\theta,\mu_{i}}(x_{i,k})\Delta t_{i,k}},x_{i,k})\big)\,.$
We note that the cost (64) corresponds to a discrete-time flow-matching, which is strongly related to the worst case orbit approximation (31) for small $\tau=\Delta t_{i,k}$. The velocity based cost (62) is an instantaneous projection error, and may be seen as a limiting case of (31) as $\lim_{\tau\rightarrow 0}$.
Theorem 5.4
If $\rho^{*}_{\theta,\mu}$ in (59) satisfies
$\rho^{*}_{\mu_{i}}$ $(x_{i,k})$ (65)
$=\arg\min_{\widetilde{A}\in\mathfrak{g}}\textnormal{dist}\big(x_{i,k+1},\Phi(e^{\widetilde{A}\Delta t_{i,k}},x_{i,k})\big)\,$,
for all $(x_{i,k},t_{i,k},\mu_{i})\in S$ then it is a non-unique solution to the optimization problem (53) with cost (59), for fixed $G,\Phi$.
Proof 5.4.1
Analogous to the proof of Theorem 5.3.
We separately optimize over $\Xi$ and $\theta$, in the following section.
5.4 Optimization Strategy
We investigate a decoupled optimization strategy, in which a first optimization identifies $(G^{*},\Phi^{*})\in\Xi$, and a second optimization identifies the minimizer $\rho^{*}_{\theta,\mu}$ of (59), (62) or (64) for fixed $(G^{*},\Phi^{*})$. The following theorem states that if the solution trajectories already lie on a group orbit, then the optimal reduced vector field recovers the FOM exactly.
Theorem 5.5 (Decoupled Optimization)
With $S$ the set of solution snapshots (18). If there exist $(G^{*},\Phi^{*})$ such that either
  1. 1.
    $x_{\mu}(t)\in\mathcal{O}(x_{\mu,0})$ for all $x_{\mu,0}\in S$.
  2. 2.
    $X_{\mu}\in\Gamma(\Delta)$.
Then there is $\rho^{*}_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$ such that $(G^{*},\Phi^{*},\rho^{*}_{\mu})$ minimizes $J$ in Eqs. (59), (62) and (64).
Proof 5.5.1
If there is $(G^{*},\Phi^{*})$ such that $x_{\mu}(t)\in\mathcal{O}(x_{\mu,0})$ for all $x_{\mu,0}\in S$, then it holds by definition for any $x_{1},x_{2}\in\mathcal{O}(x_{\mu,0})$, that there exists $g\in G$ such that
$x_{2}=\Phi(g,x_{1})\,.$ (66)
With reference to Theorem 5.4, we then construct the explicit minimizer $\rho^{*}_{\mu_{i}}(x_{i,k})=\log(g_{i,k})/\Delta t_{i,k}$ of (64). Instead, using that $X_{\mu}\in\Gamma(\Delta)$ and with reference to Theorem 5.2, the explicit solution can be constructed as $\rho_{\mu}^{*}(x)=X^{\dagger}_{\mathfrak{g}}X_{\mu}(x)$. This yields the solution to (59) and (62).
The assumptions of Theorem 5.5 directly encode that the FOM dynamics can be expressed as dynamics on $G^{*}$. The assumptions are equivalent – however, Assumption 1 can be checked more easily for discrete velocity-free measurements, and Assumption 2 can be checked more easily when the FOM dynamics are known.
If we can identify $(G^{*},\Phi^{*})\in\Xi$ for which either of the assumptions hold, the optimal $\rho^{*}_{\theta,\mu}$ is explicitly determined by Theorems 5.2,5.3 and 5.4. To this end we present a concrete algorithm for optimization in Section 7, where we perform a discrete search over select $\Xi$.
5.5 Subalgebra search
So far we did not discuss the dimension of the group $G^{*}$, which may be high-dimensional. Here, we present a means to further reduce the dimension of the ROM, assuming that a first optimizer $(G^{*},\Phi^{*},\rho_{\mu}^{*})$ was found.
We investigate the induced admissible pairs $\Xi^{G^{*},\Phi^{*}}$ (cf. Def. 5.1), which form a search space for yet lower dimensional approximations on subgroups $H^{*}\subseteq G^{*}$. To this end, we present a theorem:
Theorem 5.6 (Subalgebra search)
Given $\mathcal{M}$ and $(G^{*},\Phi^{*},\rho_{\mu}^{*})$ that minimizes the optimization problem (53) with cost (59), (62) or (64), respectively. If $\rho_{\mu}^{*}(x)\in\mathfrak{h}\subseteq\mathfrak{g}$ is fully contained in the Lie subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ corresponding to the Lie subgroup $H^{*}\subseteq G^{*}$, then $(H^{*},\Phi^{*},\rho_{\mu}^{*})$ also solves (53) for the respective cost.
By Theorem 5.6, the problem of finding a subgroup $H^{*}\subset G^{*}$ reduces to finding a subalgebra $\mathfrak{h}^{*}\subset\mathfrak{g}^{*}$ that contains $\rho_{\mu}^{*}$.
In a non-intrusive context we make a further definition:
Definition 5.7
Given $(G^{*},\Phi^{*},\rho^{*}_{\mu})$ that solves (62) or (64), and given a set of snapshots $S$ as in (61) or (63), respectively. Then we define the reduced snapshot matrix $S_{\mathfrak{g}}$ as
$S_{\mathfrak{g}}:=\{\rho^{*}_{\mu}(x_{\mu})\;|\;(x_{\mu},\mu)\in S\}\,.$ (67)
It holds that
$\rho_{\mu}^{*}(x)\in\mathfrak{h}\Leftrightarrow S_{\mathfrak{g}}\subseteq\mathfrak{h}\,.$ (68)
Theorem 5.6 and (68) allow us to identify a lower dimensional Lie group $H^{*}\subseteq G^{*}$, but this requires that $S_{\mathfrak{g}}\subseteq\mathfrak{h}$ holds for some subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$.
In practice, however, it rarely holds that $S_{\mathfrak{g}}$ is exactly contained in some subalgebra of $\mathfrak{g}$. We therefore relax the condition and approximate the dominant Lie algebra directions via PCA. Concretely, we introduce an inner product on $\mathfrak{g}$, and aim to identify a subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ by principal component analysis (PCA) of $S_{\mathfrak{g}}$ and subsequent completion of a Lie algebra $\mathfrak{h}$ by repeated application of the Lie bracket, such that “most” of $S_{\mathfrak{g}}$ is contained within $\mathfrak{h}$.
To this end, we will explore select special cases in Section 7.
6 Analytic Examples
Sections 6.1 and 6.2 highlight that MORLie is applicable to distributed systems, although the presented theory in the main article focused on the finite dimensional case.
6.1 Kolmogorov $\Xi$ width of linear transport
Consider the linear transport equation for a scalar-valued function $u\in C^{\infty}(\mathbb{R})$, and scalar $\mu_{1},\mu_{2}\in\mathbb{R}$:
$\frac{\partial{u}}{\partial{t}}+\mu_{1}\frac{\partial{u}}{\partial{x}}=0\,,\;u(x,0)=u_{\mu_{2},0}(x)=\sin(\mu_{2}x)\,.$ (69)
The solutions of (69) are $u_{\mu}(x,t)=u_{\mu_{2},0}(x-\mu_{1}t)$, providing a well understood toy-example for MOR that challenges linear subspace methods. Further consider a set of solution snapshots
$S=\{u_{\mu}(x,t)\;|\;t\in[0,T],\,(\mu_{1},\mu_{2})\in\mathcal{P}\subseteq\mathbb{R}^{2}\}\,.$ (70)
The Kolmogorov $N$-width of $S$ is well-known (cf. Sec. 2.2) to decay slowly. Instead, consider the following result:
Theorem 6.1
Let $S$ be as in (70). Further consider the Lie group $(\mathbb{R},+)$ with action $\Phi:\mathbb{R}\times C^{\infty}(\mathbb{R})\rightarrow C^{\infty}(\mathbb{R})$ given by $\Phi\big(g,u\big)(x)=f(x+g)$, and let $\Xi=\{(\mathbb{R},\Phi)\}$. Then the Kolmogorov $\Xi$-width of $S$ over the initial conditions $u_{\mu_{2},0}$ is
$d^{T}_{\Xi}(S)=0\,.$ (71)
irrespective of the choice of metric on $C^{\infty}(\mathbb{R})$.
Proof 6.1.1
We compute
$d^{T}_{\Xi}(S):=\sup_{\mu_{1},\mu_{2}\in\mathbb{R}}d_{\Xi}\big(u_{\mu}(\cdot,[0,T]),u_{\mu_{2},0}(\cdot)\big)\,.$ (72)
Expand $d_{\Xi}\big(u_{\mu}(\cdot,[0,T]),u_{\mu_{2},0}(\cdot)\big)$ to find
$d_{\Xi}\big(u_{\mu}(\cdot,[0,T]),u_{\mu_{2},0}(\cdot)\big)$ (73)
$=\inf_{\begin{subarray}{c}(G,\Phi)\in\Xi\end{subarray}}\;\sup_{u(\cdot,t)\in u_{\mu}(\cdot,[0,T])}\;\inf_{v(\cdot)\in\mathcal{O}(u_{0,\mu_{2}}(\cdot))}\textnormal{dist}(u,v)$
$=\sup_{t\in[0,T]}\;\inf_{g\in\mathbb{R}}\textnormal{dist}(u_{\mu_{2},0}(x-\mu_{1}t),u_{\mu_{2},0}(x+g))$
$=\sup_{t\in[0,T]}\textnormal{dist}(u_{\mu_{2},0}(x-\mu_{1}t),u_{\mu_{2},0}(x-\mu_{1}t))$
$=0\,.$
Where the third equality holds for $g=-\mu_{1}t$, and the fourth equality holds since $\textnormal{dist}(u,u)=0$ holds for every metric. Since no specific $\mu$ was assumed, also $d^{T}_{\Xi}(S)=0$.
Thus, the Kolmogorov $\Xi$-width of $S$ over the initial conditions is identically zero for a computationally tractable group action. This provides an example where $d^{T}_{\Xi_{N}^{\textnormal{Prop}}}(S)<d_{N}(S)$.
For completeness, a family of ROMs on the Lie group $\mathbb{R}$ can be identified by Theorem (5.2) with constant $\rho_{\mu}(u)=-\mu_{1}$, resulting in
$\dot{g}_{\mu}={R_{g}}_{*}\rho_{\mu}=-\mu_{1}\,,\;g_{\mu}(0)=0\,$, (74)
and reconstructed solutions
$\bar{u}_{\mu}(x,t)$ $=\Phi(g_{\mu}(t),u_{\mu_{2},0}(x))$ (75)
$=u_{\mu_{2},0}(x+g_{\mu}(t))$
$=u_{\mu_{2},0}(x-\mu_{1}t)$
$=u_{\mu}(x,t)\,.$
6.2 The method of freezing in MORLie
[…]
We begin by describing the method of freezing as presented in [1]. Consider the function space $U=H^{\infty}(\mathbb{R}^{n},\mathbb{R})$, an elliptic operator $\mathcal{L}_{\mu}:U\rightarrow U$, and take $u_{\mu}\in U$ subject to the following nonlinear Cauchy problem
$\dot{u}_{\mu}+\mathcal{L}_{\mu}(u_{\mu})=0\,,\;u_{\mu}(0)=u_{0}\,.$ (76)
Let there be an action $\Phi:G\times U\rightarrow U$, such that the operator $\mathcal{L}_{\mu}$ is equivariant w.r.t. $\Phi$:
$g_{\mu}^{-1}\cdot\mathcal{L}_{\mu}\big(g_{\mu}\cdot v_{\mu}\big)=\mathcal{L}_{\mu}(v_{\mu})\,$, (77)
and expand $u_{\mu}\in U$ as $g_{\mu}\in G$ acting on $v_{\mu}\in U$:
$u_{\mu}(t)=\Phi\big(g_{\mu}(t),v_{\mu}(t)\big)=:g_{\mu}(t)\cdot v_{\mu}(t)\,.$ (78)
Then the PDE (76) can be rewritten as a partial differential algebratic equation (PDAE) (79) - (81):
$\dot{v}_{\mu}$ $=-\mathcal{L}^{G}_{\mu,\widetilde{A}}\big(v_{\mu}\big)\,$, (79)
$\dot{g}_{\mu}$ $=g_{\mu}\widetilde{A}\,$, (80)
$\Psi$ $\big(\widetilde{A},\dot{v}_{\mu}\big)=0\,$, (81)
where $\mathcal{L}^{G}_{\mu,\widetilde{A}}\big(v_{\mu}\big):=\mathcal{L}_{\mu}\big(v_{\mu}\big)+X_{\widetilde{A}}(v_{\mu})$, and $\Psi\big(\widetilde{A},\dot{v}_{\mu}\big)$ is an invertible phase condition that uniquely determines $\widetilde{A}(t)\in\mathfrak{g}$.
The method of freezing proceeds as follows: high-fidelity solution snapshots of $v_{\mu}(t)$ are computed, a POD-greedy approach is used to determine a reduced basis
$\bar{U}:=\textnormal{span}(\{v_{1},\cdots,v_{N}\})\subseteq U\,$, (82)
and $v_{\mu}$ is approximated as
$v_{\mu}(t)\approx\sum_{i=1}^{N}c_{\mu}^{i}(t)v_{i}\,.$ (83)
Next, the (already discretized, high-dimensional) dynamics and phase-condition are projected to this reduced basis, using empirical operator interpolation [65] to derive $\mathbb{L}^{G}_{\mu,\widetilde{A}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ and $\mathbb{F}:\mathbb{R}^{N}\rightarrow\mathfrak{g}$ that result in a ROM of the form:
$\dot{c}_{\mu}$ $=-\mathbb{L}^{G}_{\mu,\widetilde{A}}(c_{\mu}(t))\,,\;c_{\mu}(0)=c_{\mu,0}\,$, (84)
$\dot{g}_{\mu}$ $=g_{\mu}\widetilde{A}\ \,,\;g_{\mu}(0)=e$ (85)
$\widetilde{A}(t)$ $=\mathbb{F}(c_{\mu}(t))\,.$ (86)
Finally, solutions are reconstructed as:
$u_{\mu}(t)$ $=g_{\mu}(t)\cdot v_{\mu}(t)$ (87)
$\approx\sum_{i=1}^{N}c_{\mu}^{i}(t)\big(g_{\mu}(t)\cdot v_{i})\,.$
[…]
$\mathcal{O}(\mathbb{R}^{k},G):=\bigg\{\sum_{i=1}^{N}c^{i}\big(g\cdot v_{i})\,|\,c\in\mathbb{R}^{k},g\in G\bigg\}\,.$ (88)
Lemma 6.2 (Group & Action)
Let $\widetilde{G}=G\times\mathbb{R}^{k}$ be the product of the Abelian Lie group $(\mathbb{R}^{k},+)$ and the Lie group $G$, with elements denoted by $(c_{\mu},g_{u})\in\widetilde{G}$. Further, define the augmented manifold $\mathcal{M}_{G}=G\times\mathcal{M}$ (cf. Lemma A.3) and actions $\widetilde{\Phi}_{1}:G\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$, $\widetilde{\Phi}_{2}:\mathbb{R}^{k}\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ as
$\widetilde{\Phi}_{1}\big(g,(h,u)\big)$ $=(hg,g\cdot u)\,$, (89)
$\widetilde{\Phi}_{2}\big(c,(h,u)\big)$ $=\bigg(h,u+\sum_{i=1}^{N}c^{i}(h\cdot v_{i})\bigg)\,.$ (90)
Then
  1. 1.
    The map $\widetilde{\Phi}:\widetilde{G}\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ given by
    $\widetilde{\Phi}\big((g,c),(h,u)\big)=\bigg(hg,g\cdot u+\sum_{i=1}^{N}c^{i}(h\cdot g\cdot v_{i})\bigg)\,.$ (91)
    is an action of $\widetilde{G}$ on $\mathcal{M}_{G}$.
  2. 2.
    With $\pi_{2}:\mathcal{M}_{G}\rightarrow\mathcal{M}$ the projection on the second factor, and $\mathcal{O}(\mathbb{R}^{k},G)$ as in (88), the orbit $\mathcal{O}\big((e,0)\big)$ under $\widetilde{\Phi}$ is such that
    $\pi_{2}\mathcal{O}\big((e,0)\big)=\mathcal{O}(\mathbb{R}^{k},G)\,.$ (92)
Proof 6.2.1
First, consider the actions $\Phi_{1}:G\times\mathcal{M}\rightarrow\mathcal{M}$, $\Phi_{2}:\mathbb{R}^{k}\times\mathcal{M}\rightarrow\mathcal{M}$ given by
$\Phi_{1}(g,u)$ $=g\cdot u\,$, (93)
$\Phi_{2}(c,u)$ $=u+\sum_{i=1}^{N}c^{i}v_{i}\,.$ (94)
Following Lemma A.3, commuting actions $\widetilde{\Phi}_{1}:G\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$, $\widetilde{\Phi}_{2}:\mathbb{R}^{k}\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ are constructed, and their product $\widetilde{\Phi}:\widetilde{G}\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ is guaranteed to be an action. Second, the orbit $\mathcal{O}\big((h,v)\big)$ is
$\mathcal{O}\big((h,v)\big):=\big\{\Phi\big((g,c),(h,v)\big)\,|\,(g,c)\in\widetilde{G}\big\}\,$. (95)
For $(g,\gamma_{0})=(e,0)$ this becomes
$\mathcal{O}\big((e,0)\big)=\bigg\{(g,\sum_{i=1}^{N}c^{i}(g\cdot v_{i})\big)\,|\,(g,c)\in\widetilde{G}\bigg\}\,$, (96)
and the result (92) follows by inspection.
Remark 6.3
An alternate construction of an action of $\widetilde{G}$ on the augmented manifold $M_{\mathbb{R}^{k}}$ is also possible (cf. Lemma A.3) and similarly satisfies $\pi_{2}\mathcal{O}\big((e,0)\big)=\mathcal{O}(\mathbb{R}^{k},G)$, but is not further investigated.
We are now ready to describe the method of freezing in MORLie:
Theorem 6.4 (Method of freezing in MORLie)
Given the Lie group $\widetilde{G}=G\times\mathbb{R}^{k}$, and action $\widetilde{\Phi}:\widetilde{G}\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ as in (91). Denote by $P:U\rightarrow\mathbb{R}^{k}$ the projection onto components $c^{i}$ of the basis $\{v_{1},\cdots,v_{n}\}$ spanning $\bar{U}\subseteq U$ in (82). Define $\rho:U_{G}\rightarrow\widetilde{\mathfrak{g}}$ in terms of components $\rho_{1}:U_{G}\rightarrow\mathbb{R}^{k}$ and $\rho_{2}:U_{G}\rightarrow\mathfrak{g}$:
$\rho_{1}\big(g,u)$ $=-\mathbb{L}^{G}_{\mu,\mathbb{F}(c)}(P(g^{-1}\cdot u))$ (97)
$\rho_{2}\big(g,u)$ $=\mathbb{F}(P(g^{-1}\cdot u))\,.$ (98)
Further identify $(g,u)\in\mathcal{O}\big((e,0)\big)$ with $(g,c)=(g,P(g^{-1}\cdot u))\in\widetilde{G}$. Then the MORLie ROM on $\widetilde{G}$ starting at $(g_{0},c_{0})\in\mathcal{O}\big((e,0)\big)$ reads, with $\bar{c}(t)=c_{0}+c(t)$:
$\dot{c}$ $=-\mathbb{L}^{G}_{\mu,\mathbb{F}(\bar{c})}\big(\bar{c}(t)\big)\,,\;c(0)=0\,$, (99)
$\dot{h}$ $=h\mathbb{F}\big(\bar{c}(t)\big)\,,\;h(0)=e\,.$ (100)
Let $(g_{0},u_{0})=\big(e,\,\sum_{i=1}^{N}c_{0}^{i}(v_{i})\big)$, then the reconstructed solution starting at $(g_{0},u_{0})$ is
$\widetilde{\Phi}\big((h(t),c(t)),(e,u_{0})\big)=\bigg(h(t),\,\sum_{i=1}^{N}\bar{c}^{i}(t)(h(t)\cdot v_{i})\bigg)\,.$ (101)
This solution agrees with (87) for $c_{\mu}(0)=c_{0}$ in (84) and $g_{\mu}(0)=e$ in (85), in which case $u_{\mu}(t)=\pi_{2}\widetilde{\Phi}\big((h(t),c(t)),(e,u_{0})\big)$ is the solution in the method of freezing.
Proof 6.4.1
The solutions (87) and (101) are equal if their initial conditions are equal and their dynamics are equal. This follows from uniqueness of solutions for Lipschitz dynamics. The initial conditions are $u_{0}$ in both cases. The dynamics are equal since $\dot{h}=\dot{g}_{\mu}$ and $\dot{c}=\dot{c}_{\mu}$.
Remark 6.5
Here, the key-insight is that fixed basis elements can be realized as actions of $\mathbb{R}^{k}$, and moving basis elements by the action of a product Lie group $G\times\mathbb{R}^{k}$. At this point, the formalism becomes similar to that of transformation-based methods [33, 34, 35, 36] and the shifted POD [38, 39], for which MORLie provides a geometric picture that generalizes to manifolds.
7 Numerical Examples
Here we provide basic and advanced numerical examples of the Theory in Sections 4 and 5. For a summary see Table 1. The code is available at github.com/YPWotte/MORLie.
Name Section Variables
Radial Oscillator 7.1 $\rho_{\mu}$
Rigid pointcloud 7.2 $G,\rho_{\mu}$
Sheering pointclouds 7.3 $G,\Phi,\rho_{\mu}$
Liver respiration 7.4 $\rho_{\mu}$
Table 1: Summary of degrees of freedom for optimization in numerical examples Sec. 7.1 to 7.4.
7.1 Radial oscillations
We begin with a basic example highlighting the idea of projecting a vector field to a distribution induced by a known Lie group. We consider a Riemannian manifold $\mathbb{R}^{2}$ equipped with the standard Euclidean metric. With constants $a\in\mathbb{R}$, $b\in\mathbb{N}$, $a,b\gg 1$, and parameter $\mu\in\mathbb{R}$, we define a dynamical system describing radial oscillations in polar coordinates $x_{1}=q_{1}\cos(q_{2}),x_{2}=q_{1}\sin(q_{2})$:
$\begin{cases}\dot{q}_{1}=\frac{q_{1}}{a}\sin(bq_{2})\,,\\ \dot{q_{2}}=\mu\,.\end{cases}$ (102)
We rewrite this as the FOM vector field
$X_{\mu}(q_{1},q_{2})=\frac{q_{1}}{a}\sin(bq_{2})\frac{\partial{}}{\partial{q_{1}}}+\mu\frac{\partial{}}{\partial{q_{2}}}\,.$ (103)
We further construct an action on $\mathbb{R}^{2}$ by the 1-dimensional Lie group $G=SO(2)$, expressed in polar coordinates as
$\Phi\big(\alpha,(q_{1},q_{2}))=(q_{1},q_{2}+\alpha)\,$, (104)
where the corresponding infinitesimal generator is
$X_{\dot{\alpha}}(q_{1},q_{2})=\dot{\alpha}\frac{\partial{}}{\partial{q_{2}}}\,$, (105)
which spans a distribution $\Delta=\{a(q_{1},q_{2})\frac{\partial{}}{\partial{q_{2}}}\;|\;a\in C^{\infty}(\mathbb{R}^{2})\}$. Finally, the Euclidean metric in polar coordinates reads $M=\text{d}q_{1}\otimes\text{d}q_{1}+q_{1}^{2}\text{d}q_{2}\otimes\text{d}q_{2}$, and the explicit intrusive solution in Theorem 5.2 becomes
$\rho_{\mu}^{*}(q_{1},q_{2})=X^{\dagger}\Pi_{\Delta}X_{\mu}(q_{1},q_{2})=\big(0,\mu\big)\,$, (106)
This determines the reduced dynamics $\bar{X}_{\mu}\in\Gamma(\Delta)$ as
$\bar{X}_{\mu}(q_{1},q_{2})=X_{\rho_{\mu}^{*}(q_{1},q_{2})}(q_{1},q_{2})=\mu\frac{\partial{}}{\partial{q_{2}}}\,.$ (107)
By an application of Theorem 4.1, this reduced vector field can be integrated on $SO(2)$ by solving the system
$\dot{\alpha}$ $=\rho_{\mu}^{*}\big(\Phi(\alpha,(q_{1,0},q_{2,0}))\big)=\mu\,,\;\alpha(0)=0\,.$ (108)
The resulting approximate solutions are
$\big(\bar{q_{1}}(t),\bar{q_{2}}(t)\big)=\Phi\big(\alpha(t),(q_{1,0},q_{2,0})\big)=(q_{1,0},q_{2,0}+\mu t)\,$, (109)
which are compared to FOM trajectories in Figure 2.
Figure 2: Example 7.1, MOR of the radial oscillator, showing trajectories of the full order model (FOM) and the reduced order model (ROM).
7.2 Rigid pointcloud
The following example features optimization over a set of admissible groups, and compares MORLie with a naive POD method. We will recover a rigid body motion from noisy measurements of a rigidly moving point cloud.
We consider $H(t)\in SE(3)$ with dynamics
$\dot{H}(t)=H(t)\widetilde{T}(t)\,,\;H(0)=H_{0}\,.$ (110)
for a time-dependent $\widetilde{T}(t)\in se(3)$.
With $j\in\{0,\cdots,N_{j}\}$ indexing trajectories, $i\in\{0,\cdots,N_{i}\}$ indexing particles in a given trajectory, $k\in\{0,\cdots,N_{k}\}$ indexing time-instances $t_{k}\in[0,T]$ with $t_{0}=0$, we denote the $i$-th particle on the $j$-th trajectory at time $t_{k}$ by $p^{i}_{j,k}$. Given measurement noise $\eta_{j,k}^{i}\sim\mathcal{N}(0_{3},\sigma I_{3\times 3})$, we model the position measurements by
$\begin{pmatrix}p^{i}_{j,k}\\ 1\end{pmatrix}=H(t_{k})\begin{pmatrix}p^{i}_{j,0}\\ 1\end{pmatrix}+\begin{pmatrix}\eta^{i}_{j,k}\\ 0\end{pmatrix}\,$. (111)
This represents e.g., position measurements of markers on a rigid body or landmarks on a rigid body identified from video data.
Finally, we define the system state $P_{j,k}\in\mathbb{R}^{3N}$ of $N$ particles on the $j$-th trajectory and at time $t_{k}$, by
$P_{j,k}=\begin{pmatrix}p^{1}_{j,k}\\ \vdots\\ p^{N}_{j,k}\end{pmatrix}\,$, (112)
and denote the finite set of state-time solution snapshots as $S\in\mathbb{R}^{(3N_{i}+1)\times N_{j}N_{k}}$ given by
$S:=\bigg[\begin{pmatrix}P_{0,0}\\ t_{0}\end{pmatrix},\cdots,\begin{pmatrix}P_{N_{j},N_{k}}\\ t_{N_{k}}\end{pmatrix}\bigg$]\,. (113)
7.2.1 Application of a naive POD
For a naive POD, we directly investigate the singular values of $S$ without further pre-processing. Leaving noise aside for an initial analysis:
Theorem 7.1
Assume the following:
  1. 1.
    The number of columns $N_{j}N_{k}$ and the number of rows $3N_{i}+1$ of $S$ are sufficiently large: $N_{j}N_{k}\geq 9(N_{j}+1)+3$ and $3N_{i}+1\geq 9(N_{j}+1)+3$
  2. 2.
    Each initial point cloud $P_{j,0}\in\mathbb{R}^{N_{j}}$ spans $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{p^{i}_{j,0}\;|\;i\in\{0,\cdots,N_{i}\}\}=3$
  3. 3.
    The matrices $R(t_{k})$ span $\mathbb{R}^{3\times 3}$:
    $\dim\textnormal{span}\{R(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=9$
  4. 4.
    The translations $b(t_{k})$ span $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{b(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=3$
  5. 5.
    The subspaces
    $W_{j}:=\textnormal{span}\{P_{j,k}\;|\;k\in\{0,\cdots,N_{k}\}\}$
    of $\mathbb{R}^{3N_{i}}$ associated with different point cloud trajectories $P_{j,k}$ are independent.
Then the number of non-zero singular values of the snapshot set $S$ in (113) is $9(N_{j}+1)+3$.
Proof 7.1.1
For the proof, see Appendix C.
Hence, the singular values of $S$ scale with the number of point-cloud trajectories. The assumptions for Theorem 7.1 are not strong: given randomly sampled initial pointclouds $P_{j,0}$ with sufficiently many points, they hold generically. Considering the simplicity of the underlying (6D) group dynamics (110), this is not yet a problem. If the group dynamics were more complicated, and required many different point-cloud trajectories to capture their complexity, the application of a POD without further processing of $S$ would become unfeasible.
Remark 7.2
The noise differs per particle in the point cloud, as is common in practical scenarios, and it is effectively a random vector in $\mathbb{R}^{3N_{i}}$. Expressing the noise accurately requires a POD basis that spans $\mathbb{R}^{3N_{i}}$, leading to a noise floor in the singular values.
7.2.2 Application of MORLie
As an example of MORLie, we solve (53) for the non-intrusive, velocity-free cost (64). Note that the space $\mathbb{R}^{3N_{i}}$ here constructed comes with a natural metric given by the average distance between individual particles:
$\textnormal{dist}(P_{j_{1}},P_{j_{2}})=\frac{1}{N_{i}}\sum_{i}\|p^{i}_{j_{1}}-p^{i}_{j_{2}}\|\,.$ (114)
Further, let $\textnormal{Aff}(3)=GL(3,\mathbb{R})\ltimes\mathbb{R}^{3}\subseteq GL(4,\mathbb{R})$ denote the affine group in three dimensions, i.e., $g\in\textnormal{Aff}(3)$ is identified with $A\in GL(3,\mathbb{R}),\,b\in\mathbb{R}^{3}$ and the semi-direct product $\ltimes$ denotes that the group operation is $(A_{1},b_{1})\cdot(A_{2},b_{2})=(A_{1}A_{2},b_{1}+A_{1}b_{2})$. The affine action on a single particle $p^{i}_{j,k}\in\mathbb{R}^{3}$ is a left action given by
$g\cdot p^{i}_{j,k}=Ap^{i}_{j,k}+b\,.$ (115)
We define the left action $\Phi:\textnormal{Aff}(3)\times\mathbb{R}^{3N_{i}}\rightarrow\mathbb{R}^{3N_{i}}$ by letting $g\in\textnormal{Aff}(3)$ act affinely on each particle. For the set of admissible groups we consider $\Xi^{\textnormal{Aff}(3),\Phi}$ (cf. Def. 5.1).
For fixed $(\textnormal{Aff}(3),\Phi)$, we solve for $\rho^{*}:\mathbb{R}^{3N_{i}}\rightarrow\textnormal{aff}(3)$ in two ways: first by directly applying Theorem 5.4 (in code we use a Lie-group version of the Levenberg-Marquardt algorithm for optimization, see e.g. [66, Chapter 8.4.2]), and second by approximating the FOM vector field as
$X_{\mu}(P_{j,k},t)\approx(P_{j,k}-P_{j+1,k})/\Delta t\,$, (116)
and applying Theorem 5.3 (in code the closed form solution is computed via a pseudo-inverse of $X_{\mathfrak{g}}$). These are used to construct reduced snapshot matrices $S_{\textnormal{aff}(3)}$ (cf. Def. 5.7).
The reduced snapshot matrices are then processed by Algorithm 1 (cf. Appendix B). Algorithm 1 proceeds in three steps: first, it performs a POD of $S_{\textnormal{aff}(3)}$. Singular vectors $\{\widetilde{A}_{1},\cdots,\widetilde{A}_{k}\}\subseteq\textnormal{aff}(3)$ are identified, corresponding to singular values $\sigma_{1},\cdots,\sigma_{k}$ such that $\sum^{k}_{i=1}\sigma_{i}>a\sum\sigma_{i}$, with $0<a<1$ a hyperparameter that we pick as $a=0.99$. Finally, the singular vectors are bracketed until we arrive a set $\mathfrak{g}\subseteq\textnormal{aff}(3)$ that is closed under the bracket, and is thus a Lie algebra.
For both cases of $S_{\textnormal{aff}(3)}$, the algorithm identified the 6-dimensional subalgebra $se(3)\subseteq\textnormal{aff}(3)$. Finally, we fit parameterized $\rho_{\theta}:\mathbb{R}\rightarrow se(3)$ to the identified $S_{\textnormal{aff}(3)}$, by solving
$\theta^{*}=\arg\min_{\theta}\sum_{j,k}\|\rho_{\theta}(t_{j,k})-\rho^{*}_{j,k}(P_{j,k})\|\,.$ (117)
This results in a ROM on the Lie group $SE(3)$:
$\dot{H}(t)=\rho_{\theta^{*}}(t)H(t)\,,\;H(0)=I_{4}\,$, (118)
with reconstructed solutions
$P_{i}(t)=\Phi(H(t),P_{i,0})\,.$ (119)
7.2.3 Results
We perform the computation with $N_{j}=9,N_{i}=99,N_{k}=999$. Executing the provided code on a Lenovo P15v, identification of $S_{\textnormal{aff}(3)}$ via Theorem 5.4 takes 16 seconds, and 3 seconds using the closed form Theorem 5.3. The optimization step takes 15 seconds, for an initial guess of $\rho^{*}$ being a Hermite polynomial with 100 segments, fit to every 10th entry of an average of $S_{\textnormal{aff}(3)}$ over the 10 trajectories.
Snapshots of a sample trajectory of a point-cloud are shown in Figures 3(a) - 3(c), along with the reconstructed trajectory with $S_{\textnormal{aff}(3)}$ based on Theorem 5.4. Figures 4(a) and 4(b) shows the error of the reconstructions via MORLie, and Figure 4(d) shows the error for various projections to a POD basis. Figure 4(c) shows the singular values of $S$ in (113). The singular values sharply drop at $9(N_{j}+1)+3=93$, c.f. Theorem 7.1. The singular values do not drop to zero since noise is added to each node, c.f. Remark 7.2. Figure 4(e) shows the singular values of $S_{\textnormal{aff}(3)}$ that relate to the Kolmogorov $\Xi^{\textnormal{Aff}(3)}_{N}$ width of the problem.
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
Figure 3: Example 7.2, rigidly evolving pointlcloud with noise (blue circles) and reconstructed solution (red stars).
(a) Full trajectory reconstruction errors $\textnormal{dist}\big(\bar{P}(t),P(t)\big)$ for different fits.
(b) Step ahead reconstruction error $\textnormal{dist}\big(\bar{P}_{P(t)}(\Delta t),P(t+\Delta t)\big)$ for different fits, where $\bar{P}_{P(t)}(s)$ is an approximate solution starting at $P(t)$.
(c) Singular values of $S$.
(d) Full trajectory errors $\textnormal{dist}\big(P_{POD}(t),P(t)\big)$, where $P_{POD}$ is a projection of $P$ to a POD basis.
(e) Singular values of $S_{\textnormal{aff}(3)}$.
Figure 4: Example 7.2, reconstruction errors and singular values for rigidly evolving pointcloud.
7.3 Sheering pointclouds
The following example features optimization over a set of admissible groups as well as actions. We will recover group motions of multiple particle clusters from noisy measurements of point clouds consisting of multiple affinely deforming subsets. The example is similar to the previous section, except that we consider trajectories of individual particles given by
$p^{i}_{j,k}=A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})+\eta^{i}_{j,k}\,.$ (120)
where $f(j)\in\{0,1,\cdots,n_{G}\}$ assigns to distinct affine transformations following dynamics
$\dot{g}_{f(j)}=\widetilde{A}_{f(j)}(t)g_{f(j)}\,,\;g_{f(j)}(0)=g_{f(j),0}\,.$ (121)
As a search space for groups, we consider any number of copies of the affine group, i.e., groups of the form $G=\textnormal{Aff}^{n_{G}}(3):=\textnormal{Aff}(3)\times\ldots\times\textnormal{Aff}(3)$, of a to-be-determined number $n_{G}$ of copies of $\textnormal{Aff}(3)$. For group actions, we allow any fixed assignment of particles to group clusters, which transform them by the respective affine action. This forms a large set of admissible actions $\Xi^{G,\Phi}$.
For an initial optimization inspired by Theorem 5.5, we present Algorithm 2 in Appendix B. This algorithm uses a nearest-neighbor filtering approach to split finite $S$ into clusters $S_{1},\cdots,S_{n_{G}}$, without further prior knowledge, such that points $p^{i}_{j,k}$ in $S_{f(j)}$ satisfy
$p^{i}_{j,k}\approx A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})\,.$ (122)
Thus, we arrive at $(G^{*},\Phi^{*})$, with $G^{*}$ consisting of $n_{G}$ copies of $\textnormal{Aff}(3)$, and $\Phi$ assigning affine actions to unique clusters. Then $S_{\mathfrak{g}}$ is identified as previously detailed. Again, a subalgebra is identified by Algorithm 1 resulting in a dynamics on a subgroup $(H^{*},\Phi^{*})$.
7.3.1 Results
Snapshots of a sample trajectory of two point-clouds with 100 points each are shown in Figures 5(a) - 5(c). In terms of the previous example, we have $N_{j}=9,N_{i}=199,N_{k}=999$. Executing the provided code on a Lenovo P15v, identification of $n_{G}=2$ and identification of $S_{\textnormal{aff}^{2}(3)}$ via Theorem 5.4 take 34 seconds, and 7 seconds using the closed form in Theorem 5.3. The optimization step again takes on the order of 15 seconds.
Figures 6(a) and 6(b) shows the error of the reconstructions via MORLie, Figure 6(c) shows the singular values of $S$ in (113) that relate to the Kolmogorov $N$-width of the problem, and Figure 6(d) shows the singular values of $S_{\textnormal{aff}^{2}(3)}$ that relate to the Kolmogorov $\Xi^{\textnormal{Aff}^{2}(3)}_{N}$ width of the problem.
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
Figure 5: Example 7.3, two sheering pointclouds with noise (blue circles) and reconstructed solution (red stars).
(a) Full trajectory reconstruction errors $\textnormal{dist}\big(\bar{P}(t),P(t)\big)$ for different fits.
(b) Step ahead reconstruction error $\textnormal{dist}\big(\bar{P}_{P(t)}(\Delta t),P(t+\Delta t)\big)$ for different fits, where $\bar{P}_{P(t)}(s)$ is an approximate solution starting at $P(t)$.
(c) Singular values of $S$.
(d) Singular values of $S_{\textnormal{aff}^{2}(3)}$.
Figure 6: Example 7.3, reconstruction errors and singular values for sheering pointclouds.
7.4 Tracking a liver during respiration
We present a medical application using the presented theory, highlighting practical applicability. The example concerns liver-tracking during respiration, a challenging problem that affects surgical procedures due to internal deformation and motion of the flexible liver as a patient breathes during surgery. In our example, we consider a set of points in the form of (113), collected from edge-tracking of a patient’s liver during respiration. A time-dependent trajectory on $SE(3)$ is reconstructed following the steps in Sec. 7.2 – this is shown in Figures 7(a) to 7(c). The errors of the reconstruction are shown in Figures 8(a) and 8(b).
(a) T = 0.07s.
(b) T = 58s.
(c) T = 116s.
Figure 7: Example 7.4, edge-tracking of points on a liver deforming during respiration (blue circles), and reconstructed solution assuming a motion on $SE(3)$ (red stars).
(a) Full trajectory reconstruction errors $\textnormal{dist}\big(\bar{P}(t),P(t)\big)$ for different fits.
(b) Step ahead reconstruction error $\textnormal{dist}\big(\bar{P}_{P(t)}(\Delta t),P(t+\Delta t)\big)$ for different fits, where $\bar{P}_{P(t)}(s)$ is an approximate solution starting at $P(t)$.
Figure 8: Example 7.4, reconstruction errors for example of liver deforming during respiration. For the average error $\textnormal{dist}(\bar{P},P)$ is as in (114), and for the maximum error, $\textnormal{dist}(\bar{P},P)=\max_{i}\|\bar{p}^{i}-p^{i}\|$.
7.5 Discussion
The results in Section 7.2 show that a MORLie ROM on a Lie group $SE(3)$ was successfully fit to pointcloud data undergoing a rigid motion, without assuming prior knowledge of the particular Lie group. Figure 4(e) shows that the dimension of the identified Lie group, 6, reflects in the singular values of the reduced snapshot matrix. Instead, Section 7.2.1 and Figure 4(c) show that $9(N_{j}+1)+3=93$ terms are required to capture most complexity of the data with a naive POD. Given more complicated underlying group dynamics requiring larger batches of experimental point cloud data, this is expected to make a naive POD unfeasible for data-based ROM identification – increasing the number of experiments should increase ROM accuracy, not the number of terms required.
We further note that the twist $\widetilde{T}(t)$ had to be chosen of sufficient complexity for Algorithm 1 to correctly identify the subalgebra $se(3)$. For increasing measurement noise, the Algorithm became less reliable and no longer identified $se(3)$ correctly, instead terminating at $\textnormal{aff}(3)$. Future work may investigate more robust version of Algorithm 1. Similarly, identifying $se(3)$ from Algorithm 1 still required expert knowledge, because the resulting basis elements are non-standard. This may be improved by e.g., comparing the identified sub-algebra against a library of subalgebras.
The results in Section 7.3 show that a MORLie ROM on a Lie group $\textnormal{Aff}^{2}(3)$ can be identified even when there are multiple affinely deforming and moving clusters of particles, and also when there is some overlap between the clusters. This is a first step towards more complex group-actions on clusters of particles, in which non-affine group actions may be considered.
Both results of Section 7.2 and 7.3 support that the velocity-free optimization, albeit more computationally expensive, leads to more accurate results. The authors explain this by the velocity-free optimization being more robust to the Gaussian noise $\eta$, which is amplified in the velocity-based method through the numerical differentiation (116).
Finally, the Section 7.4 shows that a rigid motion accurately models the livers motion during respiration for shallow breathing (Figures 7(a) and 7(b)), but not for deep breaths (Figure 7(c)), during which the liver significantly deforms. This reflects in the full-trajectory error in Figure 8(a), which is accurate to within 2mm for large parts of the trajectory but increases up to 7mm when deformation plays a large role. Future work may investigate nonlinear group actions to more accurately model deformations during respiration.
8 Conclusion
We presented a novel geometric framework for model order reduction via Lie groups (MORLie), giving a perspective on reduced order modeling that diverges from established linear subspace and submanifold methods for MOR. In this picture, the FOM evolves on a differentiable manifold and the ROM is cast on a Lie group acting on the FOM manifold. We defined a generalized notion of Kolmogorov $N$-widths – the group Kolmogorov $N$-width – showed how the familiar notion of Kolmogorov $N$-widths for linear subspaces is recovered as a special case, how the Kolmogorov $N$-width of the presented methods is generally lower than the linear subspace Kolmogorov $N$-width, and related it to recent nonlinear widths. We derived optimization methods for intrusive and non-intrusive optimization of the ROM and presented initial ideas for hyperreduction. In an extensive analytic examples, we apply MORLie to the transport equation and show that existing Lie group methods can be recast in the framework of MORLie. Finally, for the example of rigid and affinely evolving point-clouds we computationally implement the MOR process, and show that it scales better than linear subspace approximations, achieving a consistently low dimension.
9 Acknowledgements
We want to thank Ana Cordón Avila for sharing the data on edge-tracking in MRI scans of a patient’s liver during respiration.
Appendix A Product reductions
In Section 6.2, we combine non-commuting actions $\Phi_{G}:G\times\mathcal{M}\rightarrow\mathcal{M}$ and $\Phi_{H}:H\times\mathcal{M}\rightarrow\mathcal{M}$ into an action $\widetilde{\Phi}_{G\times H}:(G\times H)\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$. Here, we present the corresponding Theorems, alongside a further technical extension that shows how MORs via different Lie groups, i.e., $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$, can be combined into one product reduction $(G\times H,\Phi_{G\times H},\rho_{G\times H})$. We first treat the case where $\Phi_{G}$ and $\Phi_{H}$ commute.
Theorem A.1 (Product ROM for commuting actions)
Given two ROMs of $\mathcal{M},\,X\in\Gamma(T\mathcal{M})$ as $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$
$\Phi_{G}$ $:G\times\mathcal{M}\rightarrow\mathcal{M}\,,\;\rho_{G}:\mathcal{M}\rightarrow\mathfrak{g}\,$, (123)
$\Phi_{H}$ $:H\times\mathcal{M}\rightarrow\mathcal{M}\,,\;\rho_{H}:\mathcal{M}\rightarrow\mathfrak{h}\,$, (124)
such that $\Phi_{H}$ and $\Phi_{G}$ commute:
$\forall g\in G,h\in H:\;\Phi_{G,g}\circ\Phi_{H,h}=\Phi_{H,h}\circ\Phi_{G,g}\,.$ (125)
Define $\Phi_{G\times H}:G\times H\times\mathcal{M}\rightarrow\mathcal{M}$ and $\rho_{G\times H}:\mathcal{M}\rightarrow\mathfrak{g}\times\mathfrak{h}$ as
$\Phi_{G\times H}\big((g,h),x\big)$ $:=\Phi_{G}\big(g,\Phi_{H}(h,x)\big)\,$, (126)
$\rho_{G\times H}(x)$ $:=(\rho_{G}(x),\rho_{H}(x))\,.$ (127)
Then:
  1. 1.
    $\Phi_{G\times H}:G\times H\times\mathcal{M}\rightarrow\mathcal{M}$ is a group action
  2. 2.
    $(G\times H,\Phi_{G\times H},\rho_{G\times H})$ induces the approximated dynamics (cf. Def. 4.2)
    $\bar{X}(x)=X_{\rho_{G}(x)}(x)+X_{\rho_{H}(x)}(x)\,.$ (128)
Proof A.1.1 (Theorem A.1)
We begin by proving that $\Phi_{G\times H}:G\times H\times\mathcal{M}\rightarrow\mathcal{M}$ is a group action. We show that $\Phi_{G\times H}$ fulfills the homomorphism property (7):
$\Phi_{G\times H}\big((g_{1},h_{1})\cdot(g_{2},h_{2}),x\big)=\Phi_{G\times H}\big((g_{1}g_{2},h_{1}h_{2}),x\big)$
$=\Phi_{G}\big(g_{1}g_{2},\Phi_{H}\big(h_{1}h_{2},x\big)\big)$
$=\Phi_{G}\big(g_{1},\cdot\big)\circ\Phi_{G}\big(g_{2},\cdot\big)\circ\Phi_{H}\big(h_{1},\cdot\big)\circ\Phi_{H}\big(h_{2},x\big)$
$=\Phi_{G}\big(g_{1},\cdot\big)\circ\Phi_{H}\big(h_{1},\cdot\big)\circ\Phi_{G}\big(g_{2},\cdot\big)\circ\Phi_{H}\big(h_{2},x\big)$
$=\Phi_{G\times H}\big((g_{1},h_{1}),\cdot)\circ\Phi_{G\times H}\big((g_{2},h_{2}),x\big)\,.$
The first equality holds by definition of the product group, the second and third equality hold by definition (126), while the fourth equality makes use of the commutativity of $\Phi_{G}$ and $\Phi_{H}$. Finally, the fifth equality reapplies the definition (126). Next, we show that $(G\times H,\Phi_{G\times H},\rho_{G\times H})$ induces the approximated dynamics
$\bar{X}(\bar{x})=X_{\rho_{G}(\bar{x})}(\bar{x})+X_{\rho_{H}(\bar{x})}(\bar{x})\,.$ (129)
By Definition 4.2, the approximated dynamics induced by $(G\times H,\Phi_{G\times H},\rho_{G\times H})$ are given by the infinitesimal generator $X_{G\times H,\rho_{G\times H}(\bar{x})}(\bar{x})$:
$\bar{X}(\bar{x})$ $=X_{G\times H,\rho_{G\times H}(\bar{x})}(\bar{x})$
$=\frac{\text{d}}{\text{d}s}\Phi_{G\times H}\big(e^{\rho_{G\times H}(\bar{x})s},\bar{x}\big)$
$=\frac{\text{d}}{\text{d}s}\Phi_{G\times H}\big((e^{\rho_{G}(\bar{x})s},e^{\rho_{H}(\bar{x})s}),\bar{x}\big)$
$=\frac{\text{d}}{\text{d}s}\Phi_{G}\big(e^{\rho_{G}(\bar{x})s},\Phi_{H}(e^{\rho_{H}(\bar{x})s},\bar{x})\big)$
$=\frac{\text{d}}{\text{d}s}\Phi_{G}\big(e^{\rho_{G}(\bar{x})s},\bar{x}\big)+\frac{\text{d}}{\text{d}s}\Phi_{H}\big(e^{\rho_{H}(\bar{x})s},\bar{x}\big)$
$=X_{\rho_{G}(\bar{x})}(\bar{x})+X_{\rho_{H}(\bar{x})}(\bar{x})\,.$
We call the tuple $(G\times H,\Phi_{G\times H},\rho_{G\times H})$ the product ROM of $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$, and we call (128) the product approximated dynamics. A product reconstruction Theorem A.2 follows as an analog of Theorem 4.1:
Theorem A.2 (Product reconstruction)
Given a product reduction $(G\times H,\Phi_{G\times H},\rho_{G\times H})$ and the $\bar{X}(\bar{x})$ in (128), then the integral curves of
$\dot{\bar{x}}=\bar{X}(\bar{x})\,,\;\bar{x}(0)=x_{0}\,$, (130)
are given by
$\bar{x}=\Phi_{G\times H}\big((g,h)(t),x_{0}\big)\,$, (131)
where $g(t),h(t)$ satisfy (for $\Phi_{G},\Phi_{H}$ left actions):
$\dot{g}={R_{g}}_{*}\rho_{G}\big(\bar{x}(t)\big)\,,\;g(0)=e\,$, (132)
$\dot{h}={R_{h}}_{*}\rho_{H}\big(\bar{x}(t)\big)\,,\;h(0)=e\,.$ (133)
When either $\Phi_{G}$ or $\Phi_{H}$ is a right action, $R_{g}$$_{*}$ needs to be replaced by $L_{g}$$_{*}$ in (132) or $R_{h}$$_{*}$ needs to be replaced by $L_{h}$$_{*}$ in (133), respectively.
Proof A.2.1 (Theorem A.2)
To show that the reconstructed trajectory (131) is a solution of the dynamics (128), we show that its tangent vector at the point $\bar{x}\in\mathcal{M}$ is given by $\bar{X}(\bar{x})$. To this end, differentiate:
$\frac{\textnormal{d}}{\textnormal{d}t}\bar{x}(t)=$ $\,\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{G\times H}\big((g(t),h(t)),x_{0}\big)$ (134)
$=$ $\,\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{G\times H}\big((g(t),h),x_{0}\big)$
$\quad+\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{G\times H}\big((g,h(t)),x_{0}\big)$
$=$ $\,\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{G}\big(g(t),\Phi_{H}(h,x_{0})\big)$
$\quad+\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{H}\big(h(t),\Phi_{G}(g,x_{0})\big)$
$=$ $\,\frac{\textnormal{d}}{\textnormal{d}s}\Phi_{G}\big(e^{\rho_{G}(\bar{x})s},\bar{x}\big)_{|s=0}$
$\quad+\frac{\textnormal{d}}{\textnormal{d}s}\Phi_{H}\big(e^{\rho_{H}(\bar{x})s},\bar{x})_{|s=0}$
$=$ $\,X_{\rho_{G}(\bar{x})}(\bar{x})+X_{\rho_{H}(\bar{x})}(\bar{x})$
$=$ $\,\bar{X}(\bar{x})\,.$
Where the third equality uses that the group actions commute, and the fourth equality reuses that
$\frac{\textnormal{d}}{\textnormal{d}t}\Phi_{G}\big(g(t),x\big)=\frac{\textnormal{d}}{\textnormal{d}s}\Phi_{G}\big(e^{\rho_{G}(\bar{x})s},\Phi_{G}(g(t),x)\big)_{|s=0}\,.$ (135)
Given Lie groups $G,H$ and actions $\Phi_{G}:G\times\mathcal{M}\rightarrow\mathcal{M}\,,\,\Phi_{H}:H\times\mathcal{M}\rightarrow\mathcal{M}$ that do not commute, Then Theorem A.1 does not immediately apply. We make use of the following lemma:
Lemma A.3
Given any actions $\Phi_{G}:G\times\mathcal{M}\rightarrow\mathcal{M}$ and $\Phi_{H}:H\times\mathcal{M}\rightarrow\mathcal{M}$. Define the $G$-augmented manifold
$\mathcal{M}_{G}:=G\times\mathcal{M}\,.$ (136)
Then the actions $\widetilde{\Phi}_{G}:G\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ and $\widetilde{\Phi}_{H}:H\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ defined by
$\widetilde{\Phi}_{G}\big(g_{2},(g_{1},x)\big)$ $=(g_{2}g_{1},g_{2}x)\,$, (137)
$\widetilde{\Phi}_{H}\big(h,(g,x)\big)$ $=(g,ghg^{-1}x)$ (138)
commute.
Here, $\widetilde{\Phi}_{G}$ acts on the second element $\mathcal{M}$ exactly as $\Phi_{G}$ does, and $\widetilde{\Phi}_{H}$ acts on the second element of $(e,x)\in\mathcal{M}_{G}$ exactly as $\Phi_{H}$ does. This construction of commuting actions is not unique, an alternative construction uses $\mathcal{M}_{H}$ and arrives at different augmented actions.
Given this construction, a product of $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$ can be found:
Theorem A.4 (Product reduction for non-commuting actions)
Given two reductions of $\mathcal{M},\,X\in\Gamma(T\mathcal{M})$ as $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$
$\Phi_{G}$ $:G\times\mathcal{M}\rightarrow\mathcal{M}\,,\;\rho_{G}:\mathcal{M}\rightarrow\mathfrak{g}\,$, (139)
$\Phi_{H}$ $:H\times\mathcal{M}\rightarrow\mathcal{M}\,,\;\rho_{H}:\mathcal{M}\rightarrow\mathfrak{h}\,.$ (140)
Define actions $\widetilde{\Phi}_{G}:G\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ and $\widetilde{\Phi}_{H}:H\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ as in Lemma A.3, and define $\widetilde{\Phi}_{G\times H}:G\times H\times\mathcal{M}_{G}\rightarrow\mathcal{M}_{G}$ and $\rho_{G\times H}:\mathcal{M}_{G}\rightarrow\mathfrak{g}\times\mathfrak{h}$ as
$\widetilde{\Phi}_{G\times H}\big((g_{2},h),(g_{1},x)\big):=$ $\,\widetilde{\Phi}_{G}\big(g_{2},\widetilde{\Phi}_{H}(h,(g_{1},x))\big)$ (141)
$\widetilde{\rho}_{G\times H}((g,x)):=$ $\,(\rho_{G}(x),\rho_{H}(x))\,.$ (142)
Then the approximated dynamics induced by $(G\times H,\widetilde{\Phi}_{G\times H},\widetilde{\rho}_{G\times H})$ is $\bar{X}\in\Gamma(\Delta)$ with $\Delta\subseteq T\mathcal{M}_{G}$:
$\bar{X}\big((g,x)\big)=\begin{pmatrix}{R_{g}}_{*}\rho_{G}(x)\\ X_{\rho_{G}(\bar{x})}(\bar{x})+g\cdot X_{\rho_{H}(\bar{x})}(g^{-1}\cdot\bar{x})\end{pmatrix}\,$. (143)
Proof A.4.1
By Definition 4.2, the approximated dynamics induced by $(G\times H,\widetilde{\Phi}_{G\times H},\widetilde{\rho}_{G\times H})$ are given by the infinitesimal generator $X_{G\times H,\widetilde{\rho}_{G\times H}((g,\bar{x}))}((g,\bar{x}))$:
$\bar{X}\big((g,\bar{x})\big)$ $=X_{G\times H,\widetilde{\rho}_{G\times H}((g,\bar{x}))}\big((g,\bar{x})\big)$
$=\frac{\text{d}}{\text{d}s}\widetilde{\Phi}_{G\times H}\big(e^{\widetilde{\rho}_{G\times H}((g,\bar{x}))s},(g,\bar{x})\big)$
$=\frac{\text{d}}{\text{d}s}\widetilde{\Phi}_{G\times H}\big((e^{\rho_{G}(\bar{x})s},e^{\rho_{H}(\bar{x})s}),(g,\bar{x})\big)$
$=\frac{\text{d}}{\text{d}s}\widetilde{\Phi}_{G}\big(e^{\rho_{G}(\bar{x})s},\widetilde{\Phi}_{H}(e^{\rho_{H}(\bar{x})s},(g,\bar{x}))\big)$
$=\frac{\text{d}}{\text{d}s}\widetilde{\Phi}_{G}\big(e^{\rho_{G}(\bar{x})s},(g,\bar{x})\big)$
$\quad+\frac{\text{d}}{\text{d}s}\widetilde{\Phi}_{H}\big(e^{\rho_{H}(\bar{x})s},(g,\bar{x})\big)$
$=\frac{\text{d}}{\text{d}s}(e^{\rho_{G}(\bar{x})s}g,\Phi_{G}\big(e^{\rho_{G}(\bar{x})s},\bar{x})\big)$
$\quad+\frac{\text{d}}{\text{d}s}(g,g\cdot\Phi_{H}\big(e^{\rho_{H}(\bar{x})s},g^{-1}\cdot\bar{x})\big)$
$=\begin{pmatrix}{R_{g}}_{*}\rho_{G}(\bar{x})\\ X_{\rho_{G}(\bar{x})}(\bar{x})\end{pmatrix}+\begin{pmatrix}0\\ g\cdot X_{\rho_{H}(\bar{x})}(g^{-1}\cdot\bar{x})\end{pmatrix}\,$.
We call $(G\times H,\widetilde{\Phi}_{G\times H},\widetilde{\rho}_{G\times H})$ the $\mathcal{M}_{G}$-related product of $(G,\Phi_{G},\rho_{G})$ and $(H,\Phi_{H},\rho_{H})$. Also the $\mathcal{M}_{H}$-related product may be found. The differences between the two are explored in Section 6. The product reconstruction Theorem A.2 can be reused to compute integral curves on $\mathcal{M}_{G}$.
Remark A.5
The concept of a product reduction similarly induces larger group-action pairs (cf. Def. 5.1). Given group and action pairs $(G,\Phi_{G})$, $(H,\Phi_{H})$ whose actions commute, then $\Xi^{(G\times H,\Phi_{G\times H})}=\{(\widetilde{G}\times\widetilde{H},\Phi_{G\times H})\;|\;\widetilde{G}\subseteq G,\widetilde{H}\subseteq H\}$ has size $|\Xi^{(G\times H,\Phi_{G\times H})}|=|\Xi^{(G,\Phi_{G})}||\Xi^{(H,\Phi_{H})}|$. Similar results hold when the actions do not commute.
Appendix B Code
B.1 Subalgebra search
We present an algorithm that, given a finite collection $S_{\mathfrak{g}}$ of elements in a Lie algebra $\mathfrak{g}$ and an inner product on $\mathfrak{g}$, finds a subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ that contains the first $k$ singular vectors of $S_{\mathfrak{g}}$.
Algorithm 1 Subalgebra search
Input: $S_{\mathfrak{g}}$
   Output: $\mathfrak{h}\subseteq\mathfrak{g}$
[…]
5:repeat
6:  $\mathfrak{h}\leftarrow\textnormal{Basis}(\textnormal{Bracket}(\mathfrak{h},\mathfrak{h}))$
7:until $\textnormal{span}(\mathfrak{h})==\textnormal{span}(\textnormal{Bracket}(\mathfrak{h},\mathfrak{h}))$
Here, $\textnormal{Bracket}(\mathfrak{h},\mathfrak{h})$ returns $\mathfrak{h},\mathfrak{h}]\oplus\mathfrak{h}$, i.e., it uses the Lie Bracket on $\mathfrak{g}$ to bracket all elements in the basis of $\mathfrak{h}$ with each other and expands the set $\mathfrak{h}$ by the result. The algorithm is guaranteed to terminate with $\mathfrak{h}=\mathfrak{g}$, for finite $\mathfrak{g}$, but may also terminate at $\mathfrak{h}\subset\mathfrak{g}$. Accompanying code also shows how the basis of the resulting $\mathfrak{h}$, together with the Lie-bracket on $\mathfrak{g}$ can be used to construct explicitly maps such as $\exp:\mathfrak{h}\rightarrow H$, $\log:U\subseteq H\rightarrow\mathfrak{h}$ and $\text{d}\exp:T\mathfrak{h}\rightarrow\mathfrak{h}$ for numerical integration in local charts on $H$, for $G=GL(n)$. This can be more efficient than restricting the domain and co-domain of $\exp:\mathfrak{gl}(n)\rightarrow GL(n)$, $\log:U\subseteq GL(n)\rightarrow\mathfrak{gl}(n)$ and $\text{d}\exp:T\mathfrak{gl}(n)\rightarrow\mathfrak{gl}(n)$.
B.2 Group and action search
We present an algorithm that, given a collection $S$ of points $P_{j,k}\in\mathbb{R}^{3N_{i}}$ (cf. Sec. 7.3), splits them into distinct clusters and thus identifies the number of copies of the affine group $Aff(3)$ whose action can describe the motion of the points.
Algorithm 2 Clustering
Input: $S$
   Output: $n_{G}$, $S_{1},\cdots,S_{n_{G}}$
[…]
4:repeat
5:  $S_{k}\leftarrow\textnormal{RandomCluster}(S,N_{n})$
6:  $X_{k}\leftarrow\textnormal{FitInfinitesimalGenerator}(S_{k})$
7:  $S_{k}\leftarrow\textnormal{FilterByGenerator}(S,X_{k})$
[…]
9:  $k\leftarrow k+1$
10:until $|S|=0$
[…]
Here, $\textnormal{RandomCluster}(S,N_{n})$ determines a local cluster of $N_{n}$ nearest neighbors, $\textnormal{FitInfinitesimalGenerator}(S_{k})$ fits an infinitesimal generator to the velocities in $S_{k}$, and $\textnormal{FilterByGenerator}(S,X_{k})$ returns the trajectories whose velocities are described by $X_{k}$. The output of the algorithm is that the set $S$ is split into distinct groups $S_{1},\cdots,S_{n_{G}}$.
Appendix C Singular values of rigid snapshots
We prove Theorem 7.1. For convenience, we state the Theorem again:
Theorem C.1
Assume the following:
  1. 1.
    The number of columns $N_{j}N_{k}$ and the number of rows $3N_{i}+1$ of $S$ are sufficiently large: $N_{j}N_{k}\geq 9(N_{j}+1)+3$ and $3N_{i}+1\geq 9(N_{j}+1)+3$
  2. 2.
    Each initial point cloud $P_{j,0}\in\mathbb{R}^{N_{j}}$ spans $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{p^{i}_{j,0}\;|\;i\in\{0,\cdots,N_{i}\}\}=3$
  3. 3.
    The matrices $R(t_{k})$ span $\mathbb{R}^{3\times 3}$:
    $\dim\textnormal{span}\{R(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=9$
  4. 4.
    The translations $b(t_{k})$ span $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{b(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=3$
  5. 5.
    The subspaces
    $W_{j}:=\textnormal{span}\{P_{j,k}\;|\;k\in\{0,\cdots,N_{k}\}\}$
    of $\mathbb{R}^{3N_{i}}$ associated with different point cloud trajectories $P_{j,k}$ are independent.
Then the number of non-zero singular values of the snapshot set $S$ in (113) is $9(N_{j}+1)+3$.
Proof C.1.1
The number of non-zero singular values of a matrix is bounded by the number of columns $N_{j}N_{k}$, the number of rows $3N_{i}+1$, and the column rank of $S$. Given Assumption 1. that $N_{j}N_{k}$ and $3N_{i}+1$ are sufficiently large, we show that the column rank of $S$ is $9(N_{j}+1)+3$.
Each column of $S$ corresponding to trajectory $j$ at time $t_{k}$ has the form
$P_{j,k}=\begin{pmatrix}R(t_{k})p^{1}_{j,0}\\ \vdots\\ R(t_{k})p^{N_{i}}_{j,0}\end{pmatrix}+\begin{pmatrix}b(t_{k})\\ \vdots\\ b(t_{k})\end{pmatrix}$. (144)
We analyze rotational and translational contributions separately, first focusing on the rotational term.
Write as $e_{n}\in\mathbb{R}^{3}$ the $n$-th unit vector, and define vectors $\bar{P}^{(j)}_{nm}\in\mathbb{R}^{3N_{i}}$ by
$\bar{P}^{(j)}_{nm}=\begin{pmatrix}(e_{n}^{\top}p^{1}_{j,0})e_{m}\\ \vdots\\ (e_{n}^{\top}p^{N_{i}}_{j,0})e_{m}\end{pmatrix}\,$. (145)
Then
$P^{\mathrm{rot}}_{j,k}=\begin{pmatrix}R(t_{k})p^{1}_{j,0}\\ \vdots\\ R(t_{k})p^{N_{i}}_{j,0}\end{pmatrix}$ $=\sum_{n,m}R_{nm}(t_{k})\begin{pmatrix}(e_{n}^{\top}p^{1}_{j,0})e_{m}\\ \vdots\\ (e_{n}^{\top}p^{N_{i}}_{j,0})e_{m}\end{pmatrix}$ (146)
$=\sum_{n,m}R_{nm}(t_{k})\bar{P}^{(j)}_{nm}\,$. (147)
Hence the rotational part lies in the span of at most $9$ vectors. We now show that these are linearly independent. Suppose that
$\sum_{n,m=1}^{3}a_{nm}\bar{P}^{(j)}_{nm}=0.$ (148)
Then for each particle $p^{i}_{j,0}$
$\sum_{n,m}a_{nm}(e_{n}^{\top}p^{i}_{j,0})e_{m}=0\,.$ (149)
Define
$A:=\sum_{n,m}a_{nm}e_{m}e_{n}^{\top}\in\mathbb{R}^{3\times 3}.$ (150)
The previous equation is equivalent to
$Ap^{i}_{j,0}=0\quad\text{for all }i.$ (151)
By Assumption 2., the initial cloud spans $\mathbb{R}^{3}$, so this would imply $A=0$, and therefore $a_{nm}=0$ for all $n,m$. Thus, the nine vectors $\bar{P}^{(j)}_{nm}$ are linearly independent.
By Assumption 3., the rotational component for trajectory $j$ spans a $9$-dimensional subspace of $\mathbb{R}^{3N_{i}}$.
By Assumption 5., it follows that $9$ linearly independent components are required for the rotational component of each of the $N_{j}+1$ trajectories, and we arrive at $9(N_{j}+1)$ linearly independent components to express the rotational component of $S$.
By Assumption 4., the translations can be expressed by three linearly independent vectors $\bar{P}_{b,1},\cdots,\bar{P}_{b,3}$ for all trajectories, bringing the maximum number of orthogonal components to $9(N_{j}+1)+3$.
Appendix D Model order reduction on manifolds
D.1 ManiMOR
We briefly describe the recent differential geometric framework for MOR on manifolds (ManiMOR) [25].
Here, the full order model is of the form (17), along with a set of solution snapshots of the form (18). The ROM is then described as a vector field on a reduced order manifold $\mathcal{N}$ which is related to the full order manifold $\mathcal{M}$ by an embedding $\varphi:\mathcal{N}\rightarrow\mathcal{M}$.
The vector field $\bar{X}\in\mathfrak{X}(\mathcal{N})$ is given by a projection $\Pi\in C^{\infty}(T\mathcal{M},T\mathcal{N})$ of the full order vector field, and is defined as
$\bar{X}(y)=\Pi X(\varphi(y))\,$, (152)
where the projection operator $\Pi$ satisfies the projection property:
$\Pi\circ\varphi_{*}=\textnormal{id}_{T\mathcal{N}}\,.$ (153)
With respect to the main-text we note that the projection operator $\Pi_{\Delta}$ in Theorem 5.2 corresponds to $\Pi$, picking for $\mathcal{N}$ any orbit $\mathcal{N}=\mathcal{O}(x)$ and for $\varphi:\mathcal{O}(x)\rightarrow\mathcal{M}$ the canonical embedding. For more information on ManiMOR we refer to [25].
D.2 From submanifolds to distributions to Lie groups
In ManiMOR, the ROM is described on a single submanifold $\varphi(\mathcal{N})\subseteq\mathcal{M}$. Instead, we want to investigate if MOR can restrict dynamics in $\Gamma(T\mathcal{M})$ to $\Gamma(\Delta)$ with $\Delta\subseteq T\mathcal{M}$ a distribution. If $\Delta$ is regular and integrable, there will be a family of $k$-dimensional manifolds $\mathcal{N}_{x}\subseteq\mathcal{M}$ with constant $k=\dim\Delta(x)$ that contains all solutions for dynamics in $\Gamma(\Delta)$.
It is apriori unclear how to coordinatize a family of submanifolds $\mathcal{N}_{x}$, unless they are all equivalent and thus admit equivalent coordinates. Thus, we are looking for distributions $\Delta$ that induce a foliation into equivalent submanifolds $\mathcal{N}_{x}$, i.e., we are looking for a subset of regular, integrable distributions.
Finite-dimensional Lie algebras on a manifold $\mathcal{M}$ are a special case of such regular, integrable distributions. They are induced by actions of a group $G$ on the manifold $\mathcal{M}$ that are both free and proper, in which case the infinitesimal generators of the Lie algebra elements in $\mathfrak{g}$ form a Lie algebra of vector fields on $\mathcal{M}$ that spans a regular, integrable distribution. For free actions, the submanifolds $\mathcal{N}_{x}=\mathcal{O}(x)$ are all equivalent to $G$, and thus admit equivalent coordinatizations induced by $G$. This serves as a high-level motivation to study the action of Lie groups for MOR, pointing out that the decision to describe ROMs on Lie groups is still a special case from descriptions on a more general foliation into different $\mathcal{N}_{x}$.